### 17 Mar 2026 v1

# Short-horizon Duesenberry Equilibrium

Abstract.
We develop a continuous-time general equilibrium framework for economies with a heterogeneous population—modeled as a continuum—that repeatedly optimizes over short horizons under relative-income (Duesenberry-type) criteria. The cross-section evolves through a Brownian flow on a type space, transporting an effective impatience field that captures time variation in preferences induced by demographic changes, aging, and broader social shifts.
We establish three main results. First, we prove an optimal consumption–investment theorem for infinite heterogeneous populations in this Brownian-flow setting. Second, we define a short-horizon Duesenberry equilibrium—a sequential-trading (Radner-type) equilibrium in which agents repeatedly solve vanishing-horizon problems under a relative-income criterion—and provide a complete characterization and existence proof under mild regularity conditions; notably, market completeness and absence of (state-tame) arbitrage emerge endogenously from the market clearing, rather than being imposed as hypotheses. Third, we derive sharp asset-pricing implications: in equilibrium, the market price of risk is pinned down by the volatility of aggregate total wealth (financial plus human capital), implying that the equity premium is governed by the magnitudes and correlations of wealth and equity volatilities rather than by consumption volatility alone. This shifts the equity premium puzzle from an implausibly low consumption volatility to a question about the volatility of aggregate total wealth. The framework delivers explicit decompositions of the risk-free rate that are consistent with macro-finance stylized facts. All equilibrium quantities are expressed in terms of market primitives.
Key words and phrases:
Duesenberry equilibrium, Brownian flow, State-dependent utility
2020 Mathematics Subject Classification:
91B50, 91B55, 91G10, 60H10
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. Declarations of interest: none.
1. Introduction
This paper investigates intertemporal equilibrium in economies in which agents optimize over short time horizons using criteria grounded in relative income. This paper is a companion to Londoño [45], which first proposed the short-horizon optimization mechanism (see Londoño [45, Section 5.2]). The framework accommodates heterogeneous agents, including a continuum of agent types, and provides a characterization of the equilibrium along with a proof of its existence.
The proposed behavior of consumers’ decisions is based on the limiting behavior of a consumer who, over each short period, optimizes a dynamic, time-varying, isoelastic preference structure over discounted consumption and wealth (with respect to the state price process). Moreover, the dynamic preference structure may reflect changes in agents’ tastes. These changes can arise from aging or demographic shifts.
The motivation for developing a new approach to equilibrium in markets based on relative income arises from the misalignment between classical consumption theories, equilibrium models of financial markets, and empirical data. This misalignment is evidenced by several well-known puzzles, such as the “equity premium puzzle” (Mehra and Prescott [50]), the “risk-free rate puzzle” (Weil [65]), and the “risk-aversion puzzle” (Jackwerth [30]). For a discussion of the theoretical problems with the approaches to the existence of equilibrium, see Anderson and Raimondo [3]Hugonnier et al. [29]Kramkov and Predoiu [36]Riedel and Herzberg [60] and Raimondo [58].
To address these inconsistencies, Londoño [43, 45] proposed an approach for optimal consumption and investment based on optimizing a functional on consumption and wealth discounted by the state-price process. In this paper, we generalize this solution to infinite populations, incorporating population growth (see Theorem 1). A posteriori, the optimal aggregate wealth and consumption can be interpreted as a comparison of personal consumption with society’s consumption (or wealth), or with that of other consumers (see Londoño [45, Remark 17] and Remark 5).
The use of utilities that compare one’s own consumption to that of others is not new; references date back to Veblen [64] and Duesenberry [15] and include Galí [27]Abel [1]Frank [23]Frank [24]Frey and Stutzer [25]Scitovsky [62]Sen [63], and Schor [61] to cite a few.
Some works that show evidence about relative income to address problems of consumption and investment are Easterlin [16]Frey and Stutzer [25] and Easterlin et al. [17].
The solution obtained by optimizing relative wealth and consumption does not, by itself, resolve the risk-free rate puzzle or the equity premium puzzle. To address this shortcoming, Londoño [45, Section 5.2] suggested an approach that retains the optimal behavior described above, but allows consumers to solve their consumption and investment problem over short horizons, and to reset their optimization continuously as preferences and tastes change. As shown by Londoño [45, Section 5.2], this mechanism provides a satisfactory explanation for both puzzles. The present study develops and generalizes the proposed approach to a setting with infinitely many heterogeneous agents and population growth, obtaining a solution to the short-horizon optimization problem (see Theorem 2) and establishing the characterization and existence of equilibrium under weak hypotheses (see Theorem 3).
Indeed, the relative-income, short-horizon optimization mechanism proposed here breaks the tight link between aggregate consumption volatility and asset returns. By allowing preferences to move with demographics, the model produces low risk-free rates together with sizable equity premia, thereby offering a unified framework that translates the classic puzzles into a question about the volatility of aggregate total wealth—a quantity whose magnitude is empirically plausible (see Remark 8)—and provides a plausible explanation of stylized facts related to population growth, inflation, and interest rates (see Remark 10).
This work develops a framework to address state spaces characterizing finite and infinite populations of heterogeneous consumers. We assume a state space (an open set $\mathbb{D}\subset\mathbb{R}^{d}$), a Brownian flow of two-parameter processes $\varphi_{s,t}(x)$ defined on $\mathbb{D}\subset\mathbb{R}^{d}$ which models the stochastic behavior of the state variables, and provides a natural framework for modeling infinite populations. Indeed, our approach is an extension of the classical framework where the relevant variables in the economy are of the form $\int_{U}f(t,\varphi_{t}(x))\,d\mu(x)$ where $\varphi_{t}$ is an Itô process describing the evolution of state variables, $f$ is a differentiable function describing the relation between the state variables and the quantities of interest, and $\mu$ is the initial distribution on the state variables at time 0. In the context of mean field games and optimal portfolio management under competition and relative performance criteria, Lacker and Zariphopoulou [38] address the problem of mean field games for portfolio optimization with infinite populations.
We use the pricing theory based on state-tameness (see Londoño [41, 42, 45]), which does not require completeness or a priori conditions on the volatility of the price process—a feature that constitutes a major obstacle in the theory of the existence of intertemporal equilibrium (see Londoño [45] and references therein). A distinctive feature of the short-horizon Duesenberry equilibrium framework is that market completeness (in the sense of no state-tame arbitrage and well-defined pricing) emerges endogenously from the equilibrium conditions, rather than being imposed as an exogenous assumption. This contrasts sharply with the classical literature (Anderson and Raimondo [3], Hugonnier et al. [29], Kramkov and Predoiu [36], Riedel and Herzberg [60]), where “potentially dynamically complete markets” is a condition with no direct economic interpretation and must therefore be assumed a priori.
In our setting, the condition that implies completeness (the Smooth Market Condition 1) is merely a regularity requirement (continuity of $\kappa_{t}$), not a structural restriction on the market architecture. The existence of $\kappa_{t}$ satisfying $\kappa_{t}\sigma_{t}=\vartheta_{t}$, where $\sigma_{t}$ is the volatility of the market price and $\vartheta_{t}$ is the market price of risk, is guaranteed for any financial market; only its continuity is assumed. The no-arbitrage and completeness property then follows as a consequence of equilibrium, not as a prerequisite for its existence. On the other hand, for the results studied here, we only require that the income of any agent be hedgeable or insurable in a rich enough actuarial market (see Remark 3), and we do not require the hedgeability of any other financial instruments. Also, we focus on the behavior of the aggregate market price rather than on that of individual stocks. As shown by Londoño [45], Duesenberry equilibria do not imply any particular behavior for individual securities. Instead, they imply the absence of arbitrage opportunities when hedging with the market portfolio and the bond.
To model changing preferences for relative consumption and wealth over time, we define stochastic, consistent isoelastic preference structures for a population. However, we point out that most definitions and results can be extended to numerous types of utilities, including logarithmic utility functions and homogeneous preference structures (see [43, Condition 1]). We define stochastic dynamic preference structures in two steps.
The first step is to define a time-consistent state-preference structure: a family of pairs of utilities for consumption and for wealth discounted by the state-price process. Time consistency means that the same optimal consumption and investment solution is obtained across all common investment horizons (see Definition 9), allowing us to naturally address the problem of an infinitely lived agent.
The second step models the preference of each type of individual in the society (represented by a point $x\in\mathbb{D}$) as a process with values in a family of time-consistent preference structures $U_{s}$, whose realization $U_{s}(\omega)(x)$ describes the preference of the agent at time $s$ in state $\omega$. This stochastic dynamic preference structure captures the possibility that individuals can change their preferences over time, for instance, due to aging.
Some references on progressive, dynamic utilities to model changes over time of preference behavior are Musiela and Zariphopoulou [55, 56] and Karoui et al. [33], among others.
As a preliminary step toward our proposed approach, we provide an optimization result for the optimal consumption and investment (see Theorem 1), which a posteriori implies a maximization behavior on relative income (see Remark 5). Investment in markets might allow arbitrage opportunities, incompleteness, and a large heterogeneous population. Theorem 1 assumes that preferences remain unchanged as time evolves. This optimization result is a straightforward generalization to a large (heterogeneous) population of Londoño [45, Theorem 1].
As a natural step to address equilibrium, we include a second group of results consisting of Theorem 2 and Theorem 3. From these two results, we derive consumption, investment, and portfolios as the limiting outcomes of the optimal processes over time intervals in which each agent solves the optimization problem of Theorem 1. Finally, as a result of the aggregate consumption and wealth characterization, we obtain a characterization and existence theorem (Theorem 3) for the implied intertemporal equilibrium under weak conditions. Corollary 3 synthesizes the macroeconomic consequences: the state-price process factorizes as the ratio of the Duesenberry loading to aggregate income, and both the equity premium and the short interest rate admit explicit decompositions into a classical consumption-risk component and an impatience-risk component driven by demographic heterogeneity. Remark 8 shows that this decomposition provides a satisfactory quantitative resolution of the equity premium puzzle and the risk-free rate puzzle using observable data, while Remark 10 discusses qualitative implications for inflation pass-through and the secular decline of interest rates associated with demographic trends.
Also, we review the leading approaches to the equity premium puzzle—habit formation, long-run risk, and idiosyncratic-income heterogeneity—and argue that each relies on at least one latent state variable or poorly identified preference parameter to amplify the stochastic discount factor (see Remark 9).
The structure of the paper is as follows. Section 2 reviews the models and definitions presented in Londoño [43, 42, 45] in the context of markets with infinitely many agents and defines the extension of the theory of Brownian flows that is appropriate to address the problems in the paper.
In Section 3, we define consistent state preference structures and stochastic consistent preference structures; we also extend a result from Londoño [43, 44, 45] on the optimal consumption and investment theorem for utilities on state-price discounted values for infinite populations.
In Section 4, we establish Theorem 2, which characterizes the optimal consumption and investment strategy for consumers who continuously reset their optimization problem, optimizing over state-price-discounted consumption and wealth. Finally, we reach a characterization theorem on an intertemporal equilibrium where agents use the locally optimal behavior of Theorem 2 to determine their consumption and investment decisions.
In Section 5, we develop two examples that illustrate two complementary parametrizations: one treating the present value of human wealth as the primitive, the other starting from the labor-income process itself.
Finally, Section 6 presents several conclusions and possible extensions of the framework proposed in this paper. Among the most promising extensions are the application of the equilibrium framework to the term structure of interest rates (fixed-income pricing), its generalization to production economies in which the aggregate income process is endogenously determined by production technology, and the development of Duesenberry equilibria with endogenous borrowing premia, in which state-dependent household borrowing rates generate large fluctuations in total wealth without altering the aggregate pricing kernel.
Technical results on aggregation and semimartingale properties of population-weighted processes are collected in Appendix A.
2. Mathematical Framework
2.1. Stochastic Flows and Consistent Processes
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},P)$ be the canonical filtered Wiener space, where $\Omega=C([0,\infty),\mathbb{R}^{n})$, $W(t)(\omega)=\omega(t)$ is the coordinate process, $P$ is Wiener measure, $(\mathcal{F}_{t})_{t\geq 0}$ is the usual augmentation of $\sigma(W(u):0\leq u\leq t)$ by null sets, and $\mathcal{F}=\mathcal{F}_{\infty}\equiv\sigma(\cup_{t\geq 0}\mathcal{F}_{t})\cup\mathcal{N}$ where $\mathcal{N}$ is the collection of $P$-null subsets of $\sigma(\cup_{t\geq 0}\mathcal{F}_{t})$.
For each $T>0$, let $(\mathcal{F}^{T}_{s,t})=\left\{\mathcal{F}^{T}_{s,t},0\leq s\leq t\leq T\right\}$ be the two-parameter filtration where $\mathcal{F}^{T}_{s,t}$ is the smallest sub $\sigma$-field containing all null sets and $\sigma(W_{s}(u)\mid s\leq u\leq t)$, where $W_{s}(u)\equiv W(u)-W(s)$.
Assume open sets $\mathbb{D}\subset\mathbb{R}^{d}$, $\mathbb{D}^{\prime}\subset\mathbb{R}^{d^{\prime}}$ and for each $T>0$ let $\varphi_{s,t}(x,\omega)$, $0\leq s\leq t\leq T$, $x\in\mathbb{D}$, be a $\mathbb{R}^{d}$-valued continuous $C^{m,\chi}(\mathbb{D}\colon\mathbb{D}^{\prime})$-semimartingale with $0\leq\chi\leq 1$. We remind the reader that $\varphi_{s,t}(x,\omega)$ is a $C^{m,\chi}(\mathbb{D}\colon\mathbb{D}^{\prime})$ process if $\varphi_{s,t}(x,\cdot)$ is a $\mathcal{F}$ measurable continuous random field (almost everywhere with the exceptional set being in $\mathcal{F}$) taking values in $\mathbb{D}^{\prime}$ which is $m$-times continuously differentiable with respect to the spatial variable $x$. Moreover, following Kunita [37, Section 3.1], it is assumed that for $\omega$ outside the exceptional set, $\partial^{\alpha}\varphi_{s,t}(x,\omega)/\partial x^{\alpha}$ is locally Hölder continuous with exponent $\chi$. Here, $\alpha=(\alpha_{1},\cdots,\alpha_{d})$ with $\sum_{i}\alpha_{i}=m$. Additionally, we say that $\varphi_{s,t}(x)$ is a continuous semimartingale if it can be decomposed as $\varphi_{s,t}(x)\coloneqq M_{s,t}(x)+B_{s,t}(x)$. In this case, $M_{s,t}(x)$ and $B_{s,t}(x)$ are $C^{m,\chi}(\mathbb{D}\colon\mathbb{D}^{\prime})$ processes. We assume that for $s<T$, $t\to M_{s,t}(x,\cdot)$ is a continuous $(\mathcal{F}_{s,t}^{T})$ local-martingale. Additionally, $t\to\partial^{\beta}B_{s,t}(x,\omega)/\partial x^{\beta}$ is a continuous $(\mathcal{F}_{s,t}^{T})$ adapted process of bounded variation for $\beta=(\beta_{1},\cdots,\beta_{d})$ with $\sum_{i}\beta_{i}\leq m$. For $\chi=0$, this is a semimartingale of class $C^{m}$.
Next, we describe several types of consistencies. Assume that for each $s$ and $T>0$, $\psi_{s,t}(x)$, $s\leq t\leq T$, and $x\in\mathbb{D}$ is a continuous $C(\mathbb{D}\colon\mathbb{D}^{\prime})$-semimartingale (adapted to $(\mathcal{F}_{t}^{T})$). Also, $\varphi_{s,t}(x)$, $0\leq s\leq t\leq T$, $x\in\mathbb{D}$ is a continuous $C(\mathbb{D}\colon\mathbb{D})$-semimartingale. We say that the family $\psi$ is $\varphi$-consistent if for each $T>0$ there exists a set $N_{T}\in\mathcal{F}_{T}^{T}$ with $P(N_{T})=0$, such that for all $s\leq s^{\prime}\leq t\leq T$ $\omega\notin N_{T}$
$\psi_{s^{\prime},t}(\varphi_{s,s^{\prime}}(x))=\psi_{s,t}(x)\qquad\text{for all }x\in\mathbb{D}.$ (1)
In the case $\psi=\varphi$, we say $\varphi$ is consistent.
For consistent processes, we focus on cases where for all $T>0$, the mapping $\varphi_{s,t}(x)$, with $s,t\in[0,T]$, forms a Brownian flow of $C^{m}$ diffeomorphisms (for some $m\geq 0$). In particular, such flows often result from solutions to non-explosive stochastic differential equations, which we discuss next.
As examples of Brownian flows, assume that $\rho\colon\mathbb{D}\to\mathbb{R}^{d}$, and $\varrho\colon\mathbb{D}\to L(\mathbb{R}^{n}\colon\mathbb{R}^{d})$, are functions of class $C^{m,\chi}$ for an open set $\mathbb{D}$ where $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$. In this paper, $L(\mathbb{R}^{n}\colon\mathbb{R}^{d})$ stands for the set of $d\times n$ matrices. Then it is known that there exists a unique local (maximal) solution $\varphi_{s,t}(x)$ of
$d\varphi_{s,t}(x)={\rho}(\varphi_{s,t}(x))\,dt+{\varrho}(\varphi_{s,t}(x))\,dW_{s}(t),\qquad\varphi_{s,s}(x)=x$ (2)
For $x\in\mathbb{D}$, $\varrho(x)$ is a continuous function. See Kunita [37, Theorem 4.7.1, and Theorem 4.7.2] for a definition of local solutions and similar results. If $\mathbb{D}=\mathbb{R}^{d}$ and $\varrho(x)$ is uniformly Lipschitz continuous in $x$, or if the eigenvalues of $\varrho(x)\varrho^{\intercal}(x)$ are uniformly away from zero, then there exists a version of the solution $\varphi_{s,t}(x)$ which is a continuous $C^{m,\chi}(\mathbb{R}^{d}\colon\mathbb{R}^{d})$-semimartingale (see [37, Theorem 4.2.7]). A similar result can be obtained if we assume that $\varrho(x)$ is twice continuously differentiable and $\sum_{k}\partial^{2}\varrho^{i,k}(x)\varrho^{k,i}(x)/\partial x^{2}_{i}$ is bounded for all $i$ (see [37, Theorem 4.2.7]). In two of the cases mentioned above, it is possible to choose a version of $\varphi_{s,t}(x)$ that is a forward Brownian flow.
Examples of families of processes $\psi_{s,t}(x)$, $0\leq s\leq t$, and $x\in\mathbb{D}$, which are $\varphi$-consistent, are processes $\psi_{s,t}(x)=f(\varphi_{s,t}(x))$ where $f$ is some smooth or continuous function, and $\varphi$ is a consistent semimartingale process.
Moreover, in some applications, it is possible to prove that local solutions can be global (i.e., there is no explosion in finite time). Some cases of this latter type include those with an explicit solution of the SDE or those where a particular technique applies. Examples include processes studied by Londoño and Sandoval [46] or reducible SDE’s (Kloeden and Platen [34]). Most applications in finance involve solutions of SDEs on a subset of $\mathbb{R}^{d}$. For instance, most price process examples are defined on the positive cone of $\mathbb{R}^{d}$.
In this paper, we assume that the latter is the case and that $\varphi_{s,t}(x)$ is the local solution of class $C^{m,\epsilon}$ for any $0\leq\epsilon<\delta$. See Kunita [37, Theorem 4.7.1 and Theorem 4.7.2]. If for any $x\in\mathbb{D}$, $\varphi_{s,t}(x)$ has a non-explosive solution with values in $\mathbb{D}$, we will say that $\varphi_{s,t}(x)$, $x\in\mathbb{D}$ is the non-explosive solution to the stochastic differential equation on $\mathbb{D}$. Theoretical conditions for the existence of non-explosive solutions on $\mathbb{R}$ for Equation (2) are given by Kunita [37, Theorem 4.7.6]. For this latter sufficient condition, let $c(x)=\exp(2\int_{0}^{x}\rho(y)/\varrho^{2}(y))\,dy$, where $\rho(x)$, and $\varrho(x)\neq 0$ are continuous functions. It follows by [37, Theorem 4.7.6] that a sufficient condition for non-explosion on the solutions of Equation (2) is that $K(\infty)=K(-\infty)=\infty$, where $K(x)=\int_{0}^{x}2c^{-1}(z)\int_{0}^{z}c(y)/\varrho^{2}(y)\,dy\,dz$.
Without loss of generality, we can assume that $\varphi_{s,t}(x)(\omega)=\varphi_{0,t-s}(x)(\theta_{s}(\omega))$ for $\omega\in\Omega$ where $\theta\colon[0,\infty)\times\Omega\to\Omega$ is the $P$-preserving flow on $\Omega$ defined by $\theta(t,\omega)\equiv W_{\cdot}-W_{t}$. The latter approach fits in the framework of the theory of Random Dynamical Systems introduced by L. Arnold and his school (see Arnold [4]), where $(\varphi,\theta)$ is a perfect cocycle. More details can be found in Mohammed and Scheutzow [54, Theorem 2.1].
In this paper, we assume an open set $\mathbb{D}\subset\mathbb{R}^{d}$ and assume that $\varphi$ is a consistent process that, for the sake of concreteness, is assumed to be a $d$-dimensional temporally homogeneous Itô process $\varphi=(\varphi_{s,t}(x),0\leq s\leq t\leq T,x\in\mathbb{D})$, of two parameters with values in $C(\mathbb{D}\colon\mathbb{D})$. We assume that $\varphi$ is the solution of
$d\varphi^{i}_{s,t}(x)=\rho^{i}(\varphi_{s,t}(x))\,dt+\sum_{1\leq j\leq n}\varrho^{i,j}(\varphi_{s,t}(x))\,dW_{s}^{j}(t)\\ \varphi^{i}_{s,s}(x)=x_{i},i=1,\cdots,d,$ (3)
where $x=(x_{1},\cdots,x_{d})^{\intercal}$. We notice that the assumptions on $\varphi$ imply that the distributions of $\varphi_{s,t}(x)$ and $\varphi_{0,t-s}(x)$ are identical.
Assumption 1 (Local well-posedness, non-explosion, and diffeomorphic Brownian flow).
Let $\mathbb{D}\subset\mathbb{R}^{d}$ be open and assume a Brownian flow $\varphi$ on $\mathbb{D}$, which satisfies Equation (2). The drift $\rho$ and diffusion coefficient $\varrho$ are locally Lipschitz in $x$ on $\mathbb{D}$ and (jointly) continuous functions, with non-explosion / no boundary hitting in finite time. Moreover, $\rho$ and $\varrho$ are sufficiently smooth in $x$ (e.g. $C^{2,0+}$ with locally Lipschitz derivatives) so that the solution map $(s,t,x)\mapsto\varphi_{s,t}(x)$ defines a Brownian flow of $C^{2}$–diffeomorphisms on $\mathbb{D}$, as in Kunita [37].
We assume that $\rho^{i},\varrho^{i,j}$ for $j=1,\cdots,n$ and $i=1,\cdots,d$ are (jointly) continuous functions which are locally Lipschitz continuous in the spatial variable, for which global solutions to the stochastic differential equation (3) exist.
Let $\varphi_{s,t}(x)$ be a continuous, $\mathbb{D}$-valued Brownian flow of homeomorphisms as above.
Definition 1.
Let $\mathcal{M}(\mathbb{D})$ denote the set of Borel measures on $\mathbb{D}$. For $\mu\in\mathcal{M}(\mathbb{D})$ and $0\leq s\leq t$, we write $\varphi_{s,t}(\mu)(\omega)$ for the image of $\mu$ under the map $\varphi_{s,t}(\cdot)(\omega)$, namely
$\varphi_{s,t}(\mu)(F)=\mu\big(\varphi_{s,t}^{-1}(F)\big),\qquad F\in\mathcal{B}(\mathbb{D}).$
Remark 1.
Throughout the paper, we present the model with types in an open set $\mathbb{D}\subset\mathbb{R}^{d}$, following the classical formulation of stochastic flows and their induced transport of measures. This choice is only for notational convenience. All definitions and results extend verbatim when the type space is a smooth manifold $M$ and the common idiosyncratic mobility is generated by a $C^{2}$ stochastic flow of diffeomorphisms $(\varphi_{s,t})_{0\leq s\leq t}$ on $M$.
2.2. Population Dynamics and Aggregation
The semimartingale properties of population-weighted aggregates are fundamental to the equilibrium analysis. We establish these properties in Appendix A, where we prove that under suitable integrability conditions, the aggregate process $\psi^{\mu}_{s,t}$ inherits the semimartingale structure from the individual processes. The key results are:
  • Proposition 1 establishes a general Itô formula for aggregated processes;
  • Lemma 1 develops the cocycle properties of population weights that ensure consistency of aggregation across time;
  • Propositions 2 and 3 specialize to discrete and continuous populations, respectively.
In this paper, we assume two types of population structures: discrete (Assumption 3) and continuous (Assumption 4). Both cases admit social mobility, immigration, emigration, or extinction of families within types. Under either assumption, the population aggregate is a continuous semimartingale (Corollary 4).
Next, we need the concept of aggregation:
Definition 2 (Population-weighted aggregation).
Let $\varphi$ be a Brownian flow satisfying Assumption 1, and let $h:\mathbb{D}\to\mathbb{R}$ be a deterministic population growth rate, bounded from above and below, of class $C^{2,0+}$. Let $\mathcal{P}_{\varphi}\subset\mathcal{P}(\mathbb{D})$ be a family of probability measures closed under transport by $\varphi$, that is, for every $0\leq s\leq t$ and every $\nu\in\mathcal{P}_{\varphi}$,
$\varphi_{s,t}(\nu)\in\mathcal{P}_{\varphi},$
where $\varphi_{s,t}(\nu)$ denotes the image of $\nu$ under $\varphi_{s,t}$ as in Definition 1.
Define the population weight
$\Lambda_{s,t}(x)\coloneqq\exp\!\Big(\int_{s}^{t}h(\varphi_{s,u}(x))\,du\Big),\qquad 0\leq s\leq t,\;x\in\mathbb{D},$
and the measure-valued kernel $\nu_{s,t}(x,\cdot)\in\mathcal{M}(\mathbb{D})$ by
$\nu_{s,t}(x,B)\coloneqq\Lambda_{s,t}(x)\,\delta_{\varphi_{s,t}(x)}(B),\qquad B\in\mathcal{B}(\mathbb{D}).$
For any finite Borel measure $\mu$ on $\mathbb{D}$, define the action of the kernel $\nu_{s,t}$ on $\mu$ by
$(\mu\nu_{s,t})(B)\coloneqq\int_{\mathbb{D}}\nu_{s,t}(x,B)\,d\mu(x),\qquad B\in\mathcal{B}(\mathbb{D}).$
A population structure after time $s$ is a triple $\boldsymbol{\mu}=(f_{\mu},\nu_{s,t},\bar{\mu}_{s,t})$, where $f_{\mu}$ is an $\mathcal{F}_{s}$-measurable random variable such that $f_{\mu}\geq 0$ and $\mathbb{E}[|f_{\mu}|]<\infty$, and
$\mu\coloneqq f_{\mu}\,\bar{\mu}_{s,s},\qquad\bar{\mu}_{s,s}\in\mathcal{P}_{\varphi}(\mathbb{D}).$
We then define the structural no-growth composition by
$\bar{\mu}_{s,t}\coloneqq\varphi_{s,t}(\bar{\mu}_{s,s}),\qquad s\leq t,$
and the associated size-weighted population measure by
$\mu_{s,t}\coloneqq\mu\nu_{s,t},\qquad\text{that is,}\qquad\mu_{s,t}(B)=\int_{\mathbb{D}}\mathbf{1}_{\{\varphi_{s,t}(x)\in B\}}\Lambda_{s,t}(x)\,d\mu(x).$
Assume $\bar{\psi}=(\bar{\psi}_{s,t}(x))$ is a continuous $C(\mathbb{D};\mathbb{R})$ semimartingale random field. The population-weighted aggregation of $\bar{\psi}$ is the $\varphi$-consistent family defined by
$\psi^{\mu}_{s,t}\coloneqq\int_{\mathbb{D}}\bar{\psi}_{s,t}(x)\,\Lambda_{s,t}(x)\,d\mu(x)=\int_{\mathbb{D}}\psi_{s,t}(x)\,d\mu(x),$ (4)
whenever the right-hand side is a well-defined semimartingale, where
$\psi_{s,t}(x)\coloneqq\bar{\psi}_{s,t}(x)\Lambda_{s,t}(x)$
is the mass-weighted typewise total field.
Convention 1.
Throughout the paper, unless stated otherwise, we work with mass-weighted (type-wise total) fields $X_{s,t}(x)$. Their per-capita counterparts are denoted by $\bar{X}_{s,t}(x)=X_{s,t}(x)/\Lambda_{s,t}(x)$. For any type-level (per-capita) quantity $\bar{X}_{s,t}(x)$ or $\bar{X}_{t}(x)\coloneqq\bar{X}_{0,t}(x)$, we write $X_{s,t}^{\mu}$ or $X_{t}^{\mu}\coloneqq X_{0,t}^{\mu}$ for its corresponding population-weighted aggregate. A similar convention is adopted for the mass-weighted (type-wise total) fields $X_{s,t}(x)$ and $X_{t}(x)=X_{0,t}(x)$ for its corresponding population-weighted aggregate $X_{s,t}^{\mu}$ and $X_{t}^{\mu}$. This superscript notation is used only whenever a type-level counterpart exists, or when emphasizing the dependence on the population measure $\mu$; it may be omitted when the measure is clear from the context.
2.3. Financial Market Structure
In this paper, we focus on the behavior of the aggregate market price rather than on that of individual stocks. As shown by Londoño [45], Duesenberry equilibria do not imply any particular behavior for individual securities.
Since we assume that markets result from aggregated decisions, we require that the price processes and all their coefficients satisfy the fundamental consistency condition given by the following definition (for a motivation, see Corollary 4).
Definition 3.
Assume a population structure $\boldsymbol{\mu}$ with initial measure $\mu$ after time 0 (see Definition 2). For each random measure $\mu_{t}=\mu_{0,t}$, $t\geq 0$ we assume a family of (random) positive continuous $\mathcal{F}_{t+s}$-adapted semimartingale processes $X^{\mu_{t}}=(X^{\mu_{t}}_{s})$ for $0\leq s<\infty$, with the property that
$X^{\mu_{t}}_{u}=X^{\mu}_{t+u}.$ (5)
We say that the family with the above properties is a family of processes consistent with the population structure $\boldsymbol{\mu}$. We notice that a family as above is uniquely characterized by the process $X^{\mu}_{t}$ for $t\geq 0$. In this case we use interchangeably the family $X^{\boldsymbol{\mu}}$ and the process $X_{t}^{\mu}$. The meaning of the process $X_{s}^{\mu_{t}}$, for $s\geq 0$, is the evolution of process after time $t$, for $s$ units of time after, time $t$, where the aggregate of the population is given by the random measure $\mu_{t}$. Throughout this paper, we adopt the notation $X^{\mu}_{t}$, for a family of processes satisfying this definition for a population structure $\boldsymbol{\mu}$, with initial measure $\mu$ at time 0.
Next, we describe a financial market.
Definition 4 (Financial market).
Let $\boldsymbol{\mu}$ be a population structure with initial measure $\mu$ after time 0 satisfying Definition 2.
A financial market with population structure $\boldsymbol{\mu}$ is a structure $\mathfrak{M}=(P^{\mu},\boldsymbol{\mu},b^{\mu},\sigma^{\mu},\delta^{\mu},\vartheta^{\mu},r^{\mu})$ where each component is a continuous semimartingale process consistent with $\boldsymbol{\mu}$ in the sense of Definition 3, satisfying the following:
  1. (i)
    Price process. The market price process $P^{\mu}=(P^{\mu}_{t})_{t\geq 0}$ is a positive continuous semimartingale with decomposition
    $dP^{\mu}_{t}=P^{\mu}_{t}\,b^{\mu}_{t}\,dt+P^{\mu}_{t}(\sigma^{\mu}_{t})^{\intercal}\,dW(t)\qquad P^{\mu}_{0}>0,$ (6)
    where the return process $b^{\mu}=(b^{\mu}_{t})$ and the volatility matrix (column vector) process $\sigma^{\mu}=((\sigma^{\mu})^{j}_{t})_{1\leq j\leq n}$ are continuous processes consistent with $\boldsymbol{\mu}$, and $dW(t)=(dW_{1}(t)\cdots,dW_{n}(t))^{\intercal}$.
  2. (ii)
    Dividend process. The dividend rate process $\delta^{\mu}=(\delta^{\mu}_{t})$ is a continuous process consistent with $\boldsymbol{\mu}$.
  3. (iii)
    Interest rate and bond. The interest rate process $r^{\mu}=(r^{\mu}_{t})$ is a continuous process consistent with $\boldsymbol{\mu}$, and the bond price process $B^{\mu}=(B^{\mu}_{t})$ is the positive continuous semimartingale satisfying
    $dB^{\mu}_{t}=r^{\mu}_{t}\,B^{\mu}_{t}\,dt,\qquad B^{\mu}_{0}=1,\qquad t\geq 0.$ (7)
  4. (iv)
    Market price of risk. The market price of risk process $\vartheta^{\mu}=(\vartheta^{\mu}_{t})$ is a continuous $\mathbb{R}^{n}$-valued (as column vector) process consistent with $\boldsymbol{\mu}$, with $\vartheta^{\mu}_{t}\in\big(\ker(\sigma^{\mu}_{t})\big)^{\perp}$ for all $t\geq 0$, satisfying the no-arbitrage relation
    $b^{\mu}_{t}+\delta^{\mu}_{t}-r^{\mu}_{t}-\operatorname{Proj}_{\ker(\sigma^{\mu}_{t})^{\intercal}}\big(b^{\mu}_{t}+\delta^{\mu}_{t}-r^{\mu}_{t}\big)=(\sigma^{\mu}_{t})^{\intercal}\,\vartheta^{\mu}_{t}.$ (8)
Given such a financial market $\mathfrak{M}$, the state price process $H^{\mu}=(H^{\mu}_{t})$ is defined as the positive continuous semimartingale
$H^{\mu}_{t}\coloneqq\frac{Z^{\mu}_{t}}{B^{\mu}_{t}},\qquad t\geq 0,$ (9)
where
$Z^{\mu}_{t}\coloneqq\exp\left\{-\int_{0}^{t}(\vartheta^{\mu}_{u})^{\intercal}\,dW(u)-\frac{1}{2}\int_{0}^{t}\|\vartheta^{\mu}_{u}\|^{2}\,du\right\}.$ (10)
By Itô’s Lemma, the inverse state price process $(H^{\mu}_{t})^{-1}$ satisfies
$d(H^{\mu}_{t})^{-1}=(H^{\mu}_{t})^{-1}\big(r^{\mu}_{t}+\|\vartheta^{\mu}_{t}\|^{2}\big)\,dt+(H^{\mu}_{t})^{-1}(\vartheta^{\mu}_{t})^{\intercal}\,dW(t),\qquad(H^{\mu}_{0})^{-1}=1.$ (11)
Let us define $H_{s,t}^{\mu}=H^{\mu}_{t}/H^{\mu}_{s}$ for $s\leq t$. It follows that for $s\leq u\leq t$, $H^{\mu}_{u,t}H_{s,u}^{\mu}=H_{s,t}^{\mu}$ and
$d(H^{\mu}_{s,t})^{-1}=(H^{\mu}_{s,t})^{-1}\left(r^{\mu}_{t}+\|\vartheta^{\mu}_{t}\|^{2}\right)\,dt\\ +(H^{\mu}_{s,t})^{-1}(\vartheta_{t}^{\mu})^{\intercal}dW(t),\qquad(H^{\mu}_{s,s})^{-1}=1$
for $s\leq t\leq T$, so $H^{\mu}_{s,t}$ for $s\leq t$ is the state price process of the market reset at $s$.
Throughout this paper, we assume that the market satisfies Condition 1 given below, which we call the smooth market condition. We notice that since $\vartheta^{\mu}_{t}\in\ker^{\perp}(\sigma^{\mu}_{t})=\mathrm{Im}((\sigma^{\mu}_{t})^{\intercal})$ the existence of a progressive measurable function $\kappa$ with the property expressed in Equation (12) follows for any financial market (see Londoño [41]). Hence, the condition below is a weak condition on the smoothness of the aforementioned property.
Condition 1 (Smooth Market Condition).
There exists a continuous (scalar) semimartingale process $\kappa^{\mu}_{t}$ consistent with $\boldsymbol{\mu}$ in the sense of Definition 3, taking values in $\mathrm{span}((\sigma^{\mu})^{1}_{t},\cdots,(\sigma^{\mu})_{t}^{n})$, such that for all $t\geq 0$,
$\kappa^{\mu}_{t}\sigma^{\mu}_{t}=\vartheta^{\mu}_{t}.$ (12)
Remark 2.
In general, the notion of Duesenberry equilibrium implies the dynamics of the market index (the whole market) and the dynamics of the bond prices (Londoño [45, Theorem 16]). However, the Duesenberry equilibrium implies no restriction on the dynamics of individual stocks. Therefore, we assume that we can only trade bonds and the market index.
We point out that unless $\operatorname{Proj}_{\ker((\sigma^{\mu}_{t})^{\intercal})}(b^{\mu}_{t}+\delta^{\mu}_{t}-r^{\mu}_{t})=0$, there are state-tame arbitrage opportunities. Therefore, in this paper, we do not impose the absence of arbitrage at the aggregate level a priori. See Londoño [41] for definitions and characterization of state arbitrage opportunities, and state-tame portfolios.
Next, we review and extend some definitions from Londoño [43] needed to describe the Duesenberry equilibrium in the setting of this paper. These extensions are natural adaptations of the classical theory of consumption and investment (Karatzas and Shreve [32]) to the new setting proposed herein. For a detailed description of consistent and related processes, see Londoño [43] and Londoño [42].
Definition 5.
Assume a market $\mathfrak{M}$ with a population structure $\boldsymbol{\mu}$. Assume continuous, consistent with the population structure $\boldsymbol{\mu}$, real-valued random fields $\xi^{\boldsymbol{\mu}}\coloneqq(\xi^{\boldsymbol{\mu}}(x))$, $c^{\boldsymbol{\mu}}\coloneqq(c^{\boldsymbol{\mu}}(x))$, $Q^{\mu}\coloneqq(Q^{\mu})$, ${\pi}^{\boldsymbol{\mu}}\coloneqq({\pi}^{\boldsymbol{\mu}}(x))$ and $L^{\boldsymbol{\mu}}\coloneqq(L^{\boldsymbol{\mu}}(x))$, with:
Rate of consumption, Rate of endowment, Wealth, Portfolio and hedgeability:
Assume non-negative processes as above $c^{\mu}_{t}(x)$ and $Q^{\mu}_{t}(x)$ such that for all $x\in\mathbb{D}$, $E\left[\int_{0}^{\infty}H^{\mu}_{t}c^{\mu}_{t}(x)\,dt\right]<\infty$ and $E\left[\int_{0}^{\infty}H^{\mu}_{t}Q^{\mu}_{t}(x)\,dt\right]<\infty$. We say that a continuous random field $\pi_{t}^{\boldsymbol{\mu}}(x)$ is a portfolio that finances $\xi_{t}^{\mu}(x)$ with hedgeable rate of consumption and rate of endowment structure if the family of processes $\xi^{\mu}_{t}(x)$ can be financed by $\pi^{\mu}_{t}(x)$ using the endowment $Q^{\mu}_{t}(x)$ and consumption $c^{\mu}_{t}(x)$, namely:
$(B^{\mu}_{t})^{-1}\xi^{\mu}_{t}(x)=\xi^{\mu}_{0}(x)+\int_{0}^{t}(B^{\mu}_{u})^{-1}\left(Q^{\mu}_{u}(x)-c^{\mu}_{u}(x)\right)\,du+\\ \int_{0}^{t}(B^{\mu}_{u})^{-1}\pi^{\mu}_{u}(x)(\sigma^{\mu}_{u})^{\intercal}\,dW(u)+\int_{0}^{t}(B^{\mu}_{u})^{-1}\pi^{\mu}_{u}(x)(b^{\mu}_{u}+\delta^{\mu}_{u}-r^{\mu}_{u})\,du$ (13)
for all $x\in\mathbb{D}$ and $0\leq t<\infty$, and say that the family $(\xi^{\boldsymbol{\mu}},c^{\boldsymbol{\mu}},Q^{\boldsymbol{\mu}})$ as above is a wealth-consumption-and-income structure, and $\xi^{\boldsymbol{\mu}}$ is a wealth process structure, $c^{\boldsymbol{\mu}}$ is termed the rate of consumption structure, $Q^{\boldsymbol{\mu}}$ is called a rate of endowment structure. Also, $\delta^{\mu}_{t}\pi^{\mu}_{t}(x)$ is known as the return on equity.
We say that $\pi_{t}^{\mu}(x)$ is a non-arbitrage portfolio if for each $T>0$, and $0\leq t\leq T$, $H^{\mu}_{t}G^{\mu}_{t}(x)$ is uniformly bounded below, where the bound depends on each $x$ and $T$. Here $G^{\mu}_{t}(x)$ denotes the gain-in-excess process (see [42, Remark 1]) and it is defined by
$G^{\mu}_{t}(x)=B^{\mu}_{t}\int_{0}^{t}(B^{\mu}_{u})^{-1}\pi^{\mu}_{u}(x)(\sigma^{\mu}_{u})^{\intercal}\,dW(u)+\\ B^{\mu}_{t}\int_{0}^{t}(B^{\mu}_{u})^{-1}\pi^{\mu}_{u}(x)(b^{\mu}_{u}+\delta^{\mu}_{u}-r^{\mu}_{u})\,du.$
If $\pi_{t}^{\mu}(x)$ is a non-arbitrage portfolio, it follows (see Londoño [45, Remark 6]) that
$H_{t}^{\mu}\xi_{t}^{\mu}(x)+\int_{0}^{t}H_{u}^{\mu}(c_{u}^{\mu}(x)-Q^{\mu}_{u}(x))\,du=\\ \xi_{0}^{\mu}(x)+\int_{0}^{t}H^{\mu}_{u}\Big(\pi^{\mu}_{u}(x)(\sigma^{\mu}_{u})^{\intercal}-\xi^{\mu}_{u}(x)(\vartheta^{\mu}_{u})^{\intercal}\Big)\,dW(u).$ (14)
If a wealth process $(\xi^{\mu},0,Q^{\mu})$ is a hedgeable rate of consumption and rate of endowment structure, we say that $(\xi^{\mu},Q^{\mu})$ is a hedgeable wealth and rate of endowment structure.
Admissibility, and Subsistence random fields:
A subsistence random field structure $L^{\boldsymbol{\mu}}=(L^{\mu}_{t}(x))$ $x\in\mathbb{D}$ for the market $\mathfrak{M}$ is a wealth process structure where for each $T>0$, $L^{\mu}_{t}(x)H^{\mu}_{t}$, $0\leq t\leq T$ is uniformly bounded below (where the bound might depend on $x$, and $T$) such that
$E\left[H^{\mu}_{t}L^{\mu}_{t}(x)\right]<\infty$
for all $t$ and $x\in\mathbb{D}$. We say that the couple $(\pi^{\boldsymbol{\mu}},c^{\boldsymbol{\mu}})$ of portfolio on stocks structure and rate of consumption structure, is admissible for $(L^{\boldsymbol{\mu}},Q^{\boldsymbol{\mu}})$, and write $(\pi^{\boldsymbol{\mu}},c^{\boldsymbol{\mu}})\in\mathcal{A}(L^{\boldsymbol{\mu}},Q^{\boldsymbol{\mu}})$ if
$\xi^{{\mu}}_{t}(x)\geq L^{\mu}_{t}(x)\qquad\mbox{for all $0\leq t<\infty$},x\in\mathbb{D}.$
This inequality means that financial wealth cannot fall below the (possibly negative) subsistence level, interpreted as the current value of future labor income (or its insured component).
The processes $c^{\mu}_{t}=\int_{\mathbb{D}}c^{\mu}_{t}(x)\,d\mu(x)$, $Q^{\mu}_{t}=\int_{\mathbb{D}}Q^{\mu}_{t}(x)\,d\mu(x)$, $\xi^{\mu}_{t}=\int_{\mathbb{D}}\xi^{\mu}_{t}(x)\,d\mu(x)$, $\pi^{\mu}_{t}=\int_{\mathbb{D}}\pi^{\mu}_{t}(x)\,d\mu(x)$, and $L^{\mu}_{t}=\int_{\mathbb{D}}L^{\mu}_{t}(x)\,d\mu(x)$ are the aggregate rate of consumption process, aggregate rate of endowment process, aggregate wealth process, aggregate portfolio of stocks, and aggregate subsistence process, where $\mu$ denotes the population measure at time 0 introduced in Definition 2.
Next, we define a hedgeable rate of endowment structure:
Definition 6.
Assume a market $\mathfrak{M}$ with population structure $\boldsymbol{\mu}$. Assume an endowment process structure $Q^{\boldsymbol{\mu}}$, with the current value of future endowments structure $L^{\boldsymbol{\mu}}$
$L^{\mu}_{t}(x)=\frac{-1}{H^{\mu}_{t}}E\left[\int_{t}^{\infty}H^{\mu}_{u}Q^{\mu}_{u}(x)\,du\mid\mathcal{F}_{t}\right]=\\ \lim_{T\to\infty}\frac{-1}{H^{\mu}_{t}}E\left[\int_{t}^{T}H^{\mu}_{u}Q^{\mu}_{u}(x)\,du\mid\mathcal{F}_{t}\right]$
If $(L^{\boldsymbol{\mu}},Q^{\boldsymbol{\mu}})$ is a hedgeable rate of endowment structure with a portfolio $\pi^{\boldsymbol{\mu}}_{Q}$ we say that $Q^{\boldsymbol{\mu}}$ is a hedgeable rate of endowment structure. In fact, the portfolio $\pi_{Q}^{\mu}$ is indeed a non-arbitrage portfolio (see Londoño [45, Theorem 11]).
Remark 3.
In this paper, insurability of labor income refers only to the insurable (non-macro) components of labor-income risk. Specifically, we mean idiosyncratic mortality and disability risk, as well as idiosyncratic employment and wage risk, i.e. job-loss events and wage shocks that are not driven by aggregate macroeconomic factors. Macroeconomic sources of labor-income variation (aggregate productivity, business-cycle conditions, etc.) are instead captured endogenously by the equilibrium model through traded aggregate risks and the state-price process, and therefore, they are not part of the “insurable components” discussed herein.
As shown in Londoño [47], in the presence of an actuarial market satisfying life-insurance completeness, mortality, and disability risks are hedgeable through insurance contracts. Similarly, unemployment or wage-insurance contracts may hedge idiosyncratic job loss risk. In such markets, total wealth equals financial wealth plus the actuarially priced value of insured labor income.
Thus, we do not lose generality assuming hedgeability of labor income, as long as there is a sufficiently rich actuarial market in the sense of (Londoño [47]).
3. Optimal Consumption and Investment
Throughout this paper, we are mainly interested in portfolio evolution structures that we obtain as a result of the optimal behavior of consumers using isoelastic utilities:
Definition 7.
Consider a function $U\colon\left(0,\infty\right)\mapsto\mathbb{R}$ continuous, strictly increasing, strictly concave, and continuously differentiable, with $U^{\prime}(\infty)=\lim_{x\to\infty}U^{\prime}(x)=0$ and $U^{\prime}(0+)\coloneqq\lim_{x\downarrow 0}U^{\prime}(x)=\infty$. Such a function will be called a utility function.
For every utility function $U(\cdot)$, we denote by $\mathcal{I}(\cdot)$ the inverse of the derivative $U^{\prime}(\cdot)$; both of these functions are continuous, strictly decreasing and map $(0,\infty)$ onto itself with $\mathcal{I}(0+)=U^{\prime}(0+)=\lim_{x\to 0^{+}}U^{\prime}(x)=\infty$, $\mathcal{I}(\infty)=\lim_{x\to\infty}\mathcal{I}(x)=U^{\prime}(\infty)=0$. We extend $U$ by $U(0)=U(0^{+})$, and keep the same notation for the extension to $0,\infty)$ of $U$ hoping that it will be clear to the reader to which function we are referring. It is a well-known result that
$\max_{0<x<\infty}\left(U(x)-xy\right)=U(\mathcal{I}(y))-y\mathcal{I}(y),\qquad 0<y<\infty.$ (15)
Definition 8.
Consider two continuous functions $U_{1},U_{2}\colon[0,\infty)\times\left(0,\infty\right)\mapsto\mathbb{R}$, such that $U_{1}(t,\cdot)$ and $U_{2}(t,\cdot)$ are utility functions in the sense of Definition 7 for all $t\in[0,\infty)$. Moreover, if $\mathcal{I}_{1}(t,x)\coloneqq(\partial U_{1}(t,x)/\partial x)^{-1}$, and $\mathcal{I}_{2}(t,x)\coloneqq(\partial U_{2}(t,x)/\partial x)^{-1}$, where $(\partial U_{1}(t,\cdot)/\partial x)^{-1}$ and $(\partial U_{2}(t,\cdot)/\partial x)^{-1}$ denote the inverse of the derivative of $U_{1}(t,\cdot)$ and $U_{2}(t,\cdot)$ respectively, it follows that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are continuous functions. We call the aforementioned functions a state preference structure. In addition, we assume the following notations for any preference structure:
$\mathcal{X}(t,T,y)\coloneqq\mathcal{I}_{2}(T,y)+\int_{t}^{T}\mathcal{I}_{1}(s,y)\,ds$ (16)
for $0\leq t\leq T<\infty$.
We extend $U_{1}$ and $U_{2}$ by defining $U_{1}(t,0)=U_{1}(t,0^{+})$, and $U_{2}(t,0)=U_{2}(t,0^{+})$ for all $0\leq t\leq\infty$, and we keep the same notation for the extension of $U_{1}$ and $U_{2}$ to $0,\infty)\times[0,\infty)$. These utility functions are discussed in Londoño [43]. We observe that the following interpretation is natural. $U_{1}(t,\cdot)$ is the utility for consuming $c_{t}$ dollars (as seen at time 0) discounted by the state price at time $t$, and $U_{2}(t,\cdot)$ is the utility on discounted wealth by the state price (as seen at time 0) when premature death occurs at time $t$.
Natural classes of utility structures are those whose utility for terminal wealth corresponds to the utility of future consumption that is not realized because of premature death. The precise definition is as follows:
Definition 9.
A state preference structure is time consistent if for any $0\leq T^{\prime}\leq T$
$\mathcal{I}_{2}(T^{\prime},y)-\mathcal{I}_{2}(T,y)=\int_{T^{\prime}}^{T}\mathcal{I}_{1}(s,y)\,ds.$
Moreover, we say that the state preference structure has integrable inverse marginal utility if, for each $y$, $\sup_{0\leq t<\infty}\mathcal{I}_{2}(t,y)\vee\int_{0}^{\infty}\mathcal{I}_{1}(t,y)\,dt<\infty$.
If $(U_{1},U_{2})$ is a time-consistent state preference structure, the functions $\mathcal{X}(t,T,y)$ and $\mathcal{X}^{-1}(t,T,y)$ do not depend on $T$, and we denote them by $\mathcal{X}(t,y)$ and $\mathcal{X}^{-1}(t,y)$.
Remark 4.
We choose consistent isoelastic preference structures because they are scale-invariant and allow for aggregation. For each isoelastic preference structure, the choice of $U_{2}(T,y)$ implies that the solution for consumption obtained by Theorem 1 for $0\leq t\leq T^{\prime}$ with $T<T^{\prime}$ agrees on $0\leq t\leq T$, implying that $U_{2}(T,y)$, the utility on discounted final wealth $y$ can be seen as the utility for future consumption. In this way, the utility structure suggests a utility preference structure for a population $U_{1}^{\infty}(x)$ of an infinitely lived agent characterized by $x\in\mathbb{D}$, in a framework similar to that used to describe the equity premium and interest-rate puzzles.
Example 1.
Let $c>0$, $0<\alpha<1$, and $0<\beta<1$. Define $U_{1}(t,y)=ce^{-\beta t}y^{\alpha}$ and $U_{2}(t,y)=de^{-\beta t}y^{\alpha}$, where $d=c((1-\alpha)/\beta)^{1-\alpha}$. It is straightforward to see that $U_{1},U_{2}\colon[0,\infty)\times(0,\infty)\to(0,\infty)$ define a time-consistent preference structure with integrable inverse marginal utility.
Definition 10.
We say that a consistent isoelastic preference structure is a family of utility functions of the form
$U_{1}(t,y)=ce^{-\beta t}y^{\alpha},\qquad U_{2}(t,y)=c\left(\frac{1-\alpha}{\beta}\right)^{1-\alpha}e^{-\beta t}y^{\alpha},$
for $0\leq t<\infty$ with $x\in\mathbb{D}$, $0<\alpha<1$, $1>\beta>0$, and $c>0$. A preference structure for a population is defined as a parametric family of consistent isoelastic preference structures $U(x)=(U_{1}(x),U_{2}(x))$ where $U_{1}(x)=(U_{1}(t,\cdot)(x)\colon 0\leq t)$ and $U_{2}(x)=(U_{2}(t,\cdot)(x)\colon 0\leq t)$ is a state preference structure for each $x\in\mathbb{D}$. Moreover, we say that the preference structure for a population is an isoelastic preference structure if for each $x$, $U(x)$ is an isoelastic preference structure where we assume that the effective impatience rate $\gamma(x)=\beta(x)/(1-\alpha(x))$ is a smooth function of $x$, where $\beta(x)$, and $\alpha(x)$ correspond to the exponent parameter and time discount factor, respectively, for the isoelastic preference structure $(U_{1}(x),U_{2}(x))$. The interpretation is the following: the preference structure $U(x)$ represents the preference behavior toward intermediate consumption and wealth for an agent characterized by the state value $x$ for future consumption.
An isoelastic consistent stochastic preference structure for a population is a process $U_{s,t}\colon\Omega\to\mathcal{U}$ for $0\leq s\leq t$ where $\mathcal{U}$ is a class of isoelastic preference structures for a population, where the corresponding effective impatience field $\gamma_{s,t}(\omega)(x)$ of $U_{s,t}(\omega)(x)$, is $\gamma(\varphi_{s,t}(x)(\omega))$ which is a continuous random field, where $\gamma(\cdot)$ is a positive function of type $C^{2}(\mathbb{D})$. We say that the generator of the isoelastic consistent stochastic preference structure for a population is the family $U_{t}=U_{0,t}$ with effective impatience rate $\gamma(\varphi_{t}(x)(\omega))$. To avoid any unnecessary technicality, we do not assume any measurability of $U_{s,t}$ but instead, we assume continuity of $\gamma$.
A natural interpretation of an isoelastic stochastic preference structure $U_{s,t}$ is as follows: An agent is characterized by $x\in\mathbb{D}$ at time $s$, and stochastic evolution of the state value given by $\varphi_{s,t}$; Moreover, the value of utility for consumption and (final) wealth after time $t$, given that at time $s$ its state value is $x$, is the isoelastic preference structure with effective impatience rate $\gamma(\varphi_{s,t}(x)(\omega))$.
To study the optimal aggregate behavior of a population with a time-homogeneous population growth rate, we need to impose some conditions on the population structure:
Assumption 2.
$\boldsymbol{\mu}$ is a population structure with a deterministic population growth rate $h$, as in Definition 2. Assume consumers have a consistent isoelastic preference structure for a population $U(x)$ for $x\in\mathbb{D}$, where $U(x)=(U_{1}(x),U_{2}(x))$. Assume that the effective impatience rate $\gamma(x)$ is a bounded (from below and above) $C^{2}(\mathbb{D})$ function satisfying $\gamma(x)\geq\underline{\gamma}$ for some $\underline{\gamma}>0$.
We present below a theorem that describes the optimal behavior of consumers (Theorem 1). In the following theorem, which assumes no changes in tastes, we do not assume that the market is free of (state) arbitrage or (state) complete.
Theorem 1.
Assume a population structure $\boldsymbol{\mu}$ after time 0 as in Definition 2, satisfying Assumption 2. Let $\mathfrak{M}=(P,\boldsymbol{\mu},b,\sigma,\delta,\vartheta,r)$ be a financial market that satisfies the Smooth Market Condition (Condition 1), where $\kappa_{t}$ is the (continuous) process taking values in $\mathrm{span}((\sigma)^{1}_{t},\cdots,(\sigma)_{t}^{n})$ with $\kappa_{t}\sigma_{t}=\vartheta_{t}$.
Also, assume consumers with a consistent isoelastic preference structure for a population, $U(x)$ for $x\in\mathbb{D}$ with respect to $\varphi_{t}(x)$ that encodes the random dynamic of all state variables, where $U(x)=(U_{1}(x),U_{2}(x))$ and $C^{2}(\mathbb{D})$ parameter function $\gamma(x)$ where $\gamma(\varphi_{t}(x))=\gamma(x)$ is constant along the evolution of the flow. In other words, $\varphi_{t}(x)$ only encodes the population dynamics, and on each type $x$, the preference remains the same at all times. Assume a hedgeable rate of endowment (population-scaled) structure $Q_{t}(x)$ (with hedging portfolio on the stocks $\varpi_{t}(x)$ that is a non-arbitrage portfolio), and the current value of (population-scaled) future endowments structure $L_{t}(x)$, as in Definition 6.
Let $\xi$ be defined by
$\xi_{t}(x)\coloneqq L_{t}(x)+e^{-t\gamma(x)}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x))$
and $c$ defined by
$c_{t}(x)\coloneqq\gamma(x)e^{-t\gamma(x)}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x)),$
with a portfolio of stocks given by
$\pi_{t}(x)\coloneqq e^{-t\gamma(x)}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x))\kappa_{t}-\varpi_{t}(x)$
for each $t\geq 0$. Then, $(\xi_{t},{c}_{t},{Q}_{t})$ is a hedgeable cumulative consumption and endowment structure and non-arbitrage portfolio $(\pi_{t},{c}_{t})\in\mathcal{A}({L}_{t},{Q}_{t})$ that is optimal for the problem of optimal consumption and investment, where ${L}_{t}$ is the population-scaled current value of future endowments. The optimality is in the sense that for all $T>0$,
$E[\int_{0}^{T}U_{1}(t,H_{t}{c}_{t}(x))(x)\,dt+U_{2}(T,H_{T}{\xi}_{T}(x))(x)]\geq\\ E[\int_{0}^{T}U_{1}(t,H_{t}\tilde{c}_{t}(x))(x)\,dt+U_{2}(T,H_{T}\tilde{\xi}_{T}(x))(x)]$ (17)
for all $x\in\mathbb{D}$ where $({\xi}_{t},c_{t},Q_{t})$ is any other population-scaled hedgeable cumulative consumption and endowment structure.
Proof.
If $({\xi},{c},{Q})$ is defined as above, it follows that Equation (14) holds. Using Itô’s formula it follows that ${\xi}(x)$ can be financed using the endowment ${Q}_{t}(x)$ and consumption ${c}_{t}(x)$ (see Equation (13)), since $\varpi_{t}(x)$ is a non-arbitrage portfolio for $Q_{t}$. The proof follows the lines of the proof of Londoño [45, Theorem 11] and  Londoño [42, Theorem 2] with the appropriate modifications. We emphasize that the smooth market condition (Condition 1) is used in the proof of this theorem. ∎
Corollary 1.
Assume the conditions of Theorem 1 and assume Assumption 2. Assume that $(\xi,c,Q)$ is a wealth-consumption-and-income structure. We define the population-weighted effective aggregate process $\eta_{t}=\eta^{\mu}_{t}$ as follows:
$\eta_{t}\coloneqq\int_{\mathbb{D}}e^{-t\gamma(x)}\big(\xi_{0}(x)-L_{0}(x)\big)\,d\mu(x),\qquad t\geq 0.$
Then the (population-weighted) aggregate optimal wealth process satisfies
$\xi^{\mu}_{t}-L^{\mu}_{t}\coloneqq\int_{\mathbb{D}}\big(\xi_{t}(x)-L_{t}(x)\big)\,d\mu(x)=\frac{1}{H_{t}}\eta_{t},$
and
$c^{\mu}_{t}\coloneqq\int_{\mathbb{D}}c_{t}(x)\,d\mu(x)=\frac{1}{H_{t}}\left(-\partial_{t}\eta_{t}\right),$
for all $t>0$, where
$\partial_{t}\eta_{t}=\int_{\mathbb{D}}\gamma(x)\,e^{-t\gamma(x)}\big(\xi_{0}(x)-L_{0}(x)\big)\,d\mu(x),\qquad t\geq 0,$
where $L^{\mu}_{t}$ is the aggregate of the current value of future endowments $L_{t}(x)$, namely
$L^{\mu}_{t}=\int_{\mathbb{D}}L_{t}(x)\,d\mu(x).$
Moreover, the aggregate optimal portfolio in stocks is given by
$\pi_{t}^{\mu}\coloneqq\int_{\mathbb{D}}\pi_{t}(x)\,d\mu(x)=(H_{t})^{-1}\eta_{t}\,\kappa_{t}-\varpi^{\mu}_{t},$
where $\varpi^{\mu}_{t}$ is the aggregate hedging portfolio process associated with the hedging portfolio $\varpi_{t}(x)$ (of the endowment $Q_{t}(x)$), $\varpi^{\mu}_{t}=\int_{\mathbb{D}}\varpi_{t}(x)\,d\mu(x)$.
Remark 5.
The criterion in Theorem 1 offers conceptual simplicity, yet a natural objection is that individual agents cannot directly observe the financial state-price deflator $H_{t}$. This apparent difficulty dissolves once we examine the ex post structure of optimal behavior.
From Theorem 1 and Corollary 1, the comparison criterion (17) can be represented as follows: For any reference type $y\in\mathbb{D}$, define its relative total-wealth gain
$G_{t}(y)\coloneqq\frac{{\xi}_{t}(y)-{L}_{t}(y)}{{\xi}_{0}(y)-{L}_{0}(y)}.$
Then the optimization criterion used in Theorem 1 is based on the maximization of
$E\Big[\int_{0}^{T}U_{1}\big(t,H_{t}{c}_{t}(x)\big)(x)\,dt+U_{2}\big(T,H_{T}{\xi}_{T}(x)\big)(x)\Big]\\ =E\Big[\int_{0}^{T}U_{1}\Big(t,{c}_{t}(x)\frac{e^{-t\gamma(y)}}{G_{t}(y)}\Big)(x)\,dt+U_{2}\Big(T,{\xi}_{T}(x)\frac{e^{-T\gamma(y)}}{G_{T}(y)}\Big)(x)\Big]\\ =E\Big[\int_{0}^{T}U_{1}\Big(t,{c}_{t}(x)\frac{\eta_{t}}{\xi^{\mu}_{t}-L^{\mu}_{t}}\Big)(x)\,dt+U_{2}\Big(T,{\xi}_{T}(x)\frac{\eta_{T}}{\xi^{\mu}_{T}-L^{\mu}_{T}}\Big)(x)\Big],$
where $\eta_{t}\coloneqq\int_{\mathbb{D}}e^{-t\gamma(x)}({\xi}_{0}(x)-{L}_{0}(x))\,d\mu(x)$, and the parameter $\gamma$ encodes impatience.
This representation reveals a fundamental insight: ex post, agents are not maximizing wealth in the traditional sense but rather maximizing utility over relative income—specifically, their consumption and terminal wealth relative to the population-weighted total wealth $(\xi_{t}^{\mu}-L_{t}^{\mu})$. The peer-group representation (first identity) shows equivalently that agents compare themselves against the wealth growth $G_{t}(y)$ of a reference cohort. This relative income maximization—rather than absolute wealth maximization—is the defining characteristic that justifies calling our framework a Duesenberry equilibrium, connecting it directly to Duesenberry [15]’s relative income hypothesis.
From a practical perspective, a typical agent does not need to observe $H_{t}$ directly. Instead, by tracking the wealth dynamics of a reference peer group—information that is plausibly available through ordinary social and economic observation—the agent obtains sufficient information to make consumption and investment decisions over short horizons. The ratio $G_{t}(y)$ summarizes the growth of total wealth (financial plus human) for the reference cohort and delivers an operational proxy for $H_{t}$ through relative comparisons. The last identity confirms that, in equilibrium, this peer-group proxy coincides with the aggregate benchmark.
It is important to emphasize that this relative comparison criterion operates only for short-horizon decisions. Agents use peer comparisons to determine their optimal consumption and investment over each short interval; the full dynamic solution emerges by rolling forward these local decisions, as we explain in Definition 11, Theorem 2, and Theorem 3. Beyond the short-horizon optimization step, the interpretation of “keeping up with the Joneses” no longer applies—the rolling procedure is purely mechanical.
Remark 6.
A potential source of confusion is that decisions are made by individual households (or families), not by population aggregates. The point is that under CRRA (isoelastic) preferences and infinite divisibility, a household with time-varying mass can be represented by a type block whose natural quantities are mass-scaled.
For each type $x$, we consider per-capita wealth, consumption, and endowment $\bar{\xi}_{t}(x):=\Lambda^{-1}_{t}(x)\xi_{t}(x)$, $\bar{c}_{t}(x):=\Lambda^{-1}_{t}(x)c_{t}(x)$, $\bar{Q}_{t}(x):=\Lambda^{-1}_{t}(x)Q_{t}(x)$, $\bar{\pi}_{t}(x):=\Lambda^{-1}_{t}(x)\pi_{t}(x)$. Population growth changes $\Lambda_{t}(x)$ by introducing new mass (new entrants) and hence new aggregate resources; it is therefore not economically meaningful to require a per-capita self-financing identity for $(\bar{\xi}_{t}(x),\bar{c}_{t}(x),\bar{Q}_{t}(x))$, since this would implicitly treat the inflow due to mass growth as a cash flow replicable by trading. Feasibility/admissibility is instead imposed on the block-level processes $(\xi,c,Q)$, for which the standard deflated budget identity holds under the market deflator $H$.
Equivalently, one may view the model through a short-horizon (rolling) lens: over a small interval $t,t+\Delta]$ the mass is approximately constant, so each household solves a standard consumption–investment problem; the update from $t$ to $t+\Delta$, then incorporates the net entry of mass at rate $h$, and the continuous-time limit yields precisely the block-level feasibility used in Theorem 1.
4. Short-horizon Duesenberry Equilibrium: Characterization and Existence
Next, we describe the setting of the model we propose for the equilibrium.
Definition 11.
Assume a financial market $\mathfrak{M}=(P,\boldsymbol{\mu},b,\sigma,\delta,\vartheta,r)$ that satisfies the Smooth Market Condition (Condition 1). Suppose an isoelastic consistent preference structure for a population $U$ with corresponding parameter $\gamma(\varphi_{t}(x)(\omega))$ of $U_{t}(\omega)(x)$, where $\gamma(\cdot)$ is a function of type $C^{2}(\mathbb{D})$, satisfying Assumption 2. Assume an aggregate hedgeable (by state-tame portfolios) rate of consumption and rate of endowment structures $(\xi,c,Q)$ and non-arbitrage portfolio $\pi$. Let $\Gamma=\{s_{0}=0<s_{1}<\cdots<s_{m}=T\}$ be a partition of the interval $[0,T]$. The endowment and consumption structure associated with $\Gamma$ and preference structure $U$ satisfying Assumption 2, are the endowment and consumption processes defined by the wealth-consumption-and-income structure $(\xi^{\Gamma},c^{\Gamma},Q)$ explained below where $\xi^{\Gamma}=(\xi^{\Gamma}_{t}(x))$ is a continuous wealth process and $c^{\Gamma}=(c^{\Gamma}_{t}(x))$ is the piecewise continuous process, where on each interval $s_{i},s_{i+1}]$ they are the optimal processes described in Theorem 1. In other words, for any $x\in\mathbb{D}$, $s_{k}\leq t<s_{k+1}$ and we denote $s_{k}^{+}=t$ and $s_{i}^{+}=s_{i+1}$ for $0\leq i\leq k-1$:
$\xi^{\Gamma}_{t}(x)-L_{t}(x)\coloneqq e^{-(t-s_{k})\gamma(\varphi_{s_{k}}(x))}H_{s_{k},t}^{-1}(\xi^{\Gamma}_{s_{k}}(x)-L_{s_{k}}(x))=\\ (\xi_{0}(x)-L_{0}(x))\prod_{i=0}^{i=k}e^{-(s_{i}^{+}-s_{i})\gamma(\varphi_{{s_{i}}}(x))}H_{s_{i},s_{i}^{+}}^{-1},$ (18)
and $c^{\Gamma}$ is given by
$c^{\Gamma}_{t}(x)\coloneqq\gamma(\varphi_{s_{k}}(x))e^{-\gamma(\varphi_{s_{k}}(x))(t-s_{k})}H_{s_{k},t}^{-1}(\xi^{\Gamma}_{s_{k}}(x)-L_{s_{k}}(x))\\ =\gamma(\varphi_{s_{k}}(x))\left(\xi^{\Gamma}_{t}(x)-L_{t}(x)\right),$ (19)
and non-arbitrage portfolio $\pi^{\Gamma}_{t}=(\pi^{\Gamma}_{t}(x))$ defined by:
$\pi^{\Gamma}_{t}(x)\coloneqq e^{-(t-s_{k})\gamma(\varphi_{s_{k}}(x))}H_{s_{k},t}^{-1}(\xi^{\Gamma}_{s_{k}}(x)-L_{s_{k}}(x))\kappa_{t}-\varpi_{t}(x)\\ =\left(\xi^{\Gamma}_{t}(x)-L_{t}(x)\right)\kappa_{t}-\varpi_{t}(x),$ (20)
where $\varpi_{t}(x)$ is the portfolio that hedges the rate of endowment structure $Q$, and $L_{t}(x)$ is the current value of the future endowment structure.
Passing to the limit as $\|\Gamma_{n}\|\to 0$ in (18)–(20) yields the following theorem.
Theorem 2.
Assume a financial market $\mathfrak{M}=(P,\boldsymbol{\mu},b,\sigma,\delta,\vartheta,r)$ that satisfies the Smooth Market Condition (Condition 1). Assume an isoelastic consistent preference structure for a population $U$ with effective impatience field $\gamma(\varphi_{t}(x)(\omega))$ of $U_{t}(\omega)(x)$, where $\gamma(\cdot)$ is a function of type $C^{2}(\mathbb{D})$, satisfying Assumption 2. Let $\xi$ be the family of continuous semimartingales
$\xi_{t}(x)\coloneqq L_{t}(x)+e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x))$
and $c$ the family of continuous semimartingales given by
$c_{t}(x)\coloneqq\gamma(\varphi_{t}(x))e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x)),$
and non-arbitrage portfolio on stocks $\pi$ with generator $\pi_{t}(x)$ given by
$e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}H_{t}^{-1}(\xi_{0}(x)-L_{0}(x))\kappa_{t}-\varpi_{t}(x)$
for any $x\in\mathbb{D}$, where $\varpi_{t}(x)$ is the portfolio that hedges the rate of endowment structure $Q$. Then, $(\xi,c,Q)$ is a hedgeable cumulative consumption and endowment structure, and $(\pi,c)\in\mathcal{A}(L,Q)$ is a non-arbitrage portfolio that arises as the pointwise limit of the sequence $(\xi^{\Gamma_{n}},c^{\Gamma_{n}},Q)$ associated with partitions $\Gamma_{n}$, where each $(\xi^{\Gamma_{n}},c^{\Gamma_{n}},Q)$ is optimal on each subinterval of $\Gamma_{n}$ in the sense of Theorem 1 (cf. Definition 11), and $L_{t}(x)$ is the current value of future endowments for each type $x\in\mathbb{D}$.
Proof.
Fix $x\in\mathbb{D}$ and $T>0$. For any partition $\Gamma=\{0=s_{0}<\cdots<s_{m}=T\}$, define $\xi^{\Gamma}$$c^{\Gamma}$, and $\pi^{\Gamma}$ by rolling forward the optimal policy of Theorem 1 on each interval $s_{i},s_{i+1}]$ with frozen impatience $\gamma(\varphi_{s_{i}}(x))$ and initial wealth $\xi^{\Gamma}_{s_{i}}(x)$. This yields the explicit recursions (18)–(20), and in particular $\big(\xi^{\Gamma},c^{\Gamma},Q\big)$ is hedgeable and $\big(\pi^{\Gamma},c^{\Gamma}\big)$ is a state-tame (non-arbitrage) portfolio for every $\Gamma$.
Using the cocycle properties $H_{s,t}=H_{t}/H_{s}$ (and the flow/weight cocycle, cf. Appendix A), the product form in (18) can be written as
$\xi_{t}^{\Gamma}(x)-L_{t}(x)=H_{t}^{-1}\,(\xi_{0}(x)-L_{0}(x))\,\exp\!\Big(-\sum_{i:\,s_{i}<t}\gamma(\varphi_{s_{i}}(x))(s_{i+1}\wedge t-s_{i})\Big).$
Since $u\mapsto\gamma(\varphi_{u}(x))$ is continuous, the Riemann sums converge as $\|\Gamma\|\to 0$ to $\int_{0}^{t}\gamma(\varphi_{u}(x))\,du$, hence $\xi_{t}^{\Gamma}(x)\to\xi_{t}(x)$ pointwise for all $t\in[0,T]$, with $\xi$ as stated in the theorem. Similarly, if $c^{\Gamma}$, and $\pi^{\Gamma}$ are the piecewise continuous processes defined by Equation (19) and Equation (20) respectively, then $c^{\Gamma}_{t}(x)\to c_{t}(x)$, and $\pi^{\Gamma}_{t}(x)\to\pi_{t}(x)$ as $\|\Gamma\|\to 0$ by continuity of $\gamma(\varphi_{\cdot}(x))$, $\kappa$, and $\varpi$. By Assumption 2, it follows that $c_{t}(x)$ is a rate of consumption process, satisfying the integrability condition $\mathbb{E}\!\left[\int_{0}^{\infty}H_{t}\,c_{t}(x)\,dt\right]<\infty$ for all $x\in\mathbb{D}$.
Moreover, since the triple $(\xi^{\Gamma},c^{\Gamma},Q)$ satisfies the financeability Equation (13), and since Assumption 2 holds, the stochastic dominated convergence theorem holds, and a localization argument yields that $(\xi,c,Q)$ satisfies (13) and is therefore a wealth, consumption, and income structure.
We also notice that $\pi_{t}(x)$ is a non-arbitrage portfolio since the original portfolio on each subinterval is a state-tame non-arbitrage portfolio (see Londoño [45, Remark 6]).
Finally, since $\xi_{t}^{\Gamma}(x)\geq L_{t}(x)$ on each subinterval, it follows that $\xi_{t}(x)\geq L_{t}(x)$, implying that $(\pi,c)$ is admissible for $(L,Q)$. ∎
Corollary 2 (Local optimal aggregate).
Assume the conditions of Theorem 2 with optimal solution $(\xi,c,Q)$, where Assumption 2 holds. Consider the aggregate process $\eta_{t}=\eta^{\mu}_{t}$ associated with the field defined by
$\eta_{t}\coloneqq\int_{\mathbb{D}}e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}y(x)\,d\mu(x),\qquad t\geq 0,$ (21)
where $\gamma(x)$ is the effective impatience rate from the preference structure, and $y(x)\coloneqq\xi_{0}(x)-L_{0}(x)$ is the initial total wealth of type $x$. Then the population-weighted aggregate of the optimal wealth processes satisfies
$\xi^{\mu}_{t}-L^{\mu}_{t}\coloneqq\int_{\mathbb{D}}\big(\xi_{t}(x)-L_{t}(x)\big)\,d\mu(x)=\frac{1}{H_{t}}\eta_{t},$ (22)
where
$L^{\mu}_{t}\coloneqq\int_{\mathbb{D}}L_{t}(x)\,d\mu(x)$
is the current value of future endowments $L_{t}(x)$.
The population-weighted aggregate of the optimal consumption process is given in terms of $\eta$ by
$c^{\mu}_{t}\coloneqq\int_{\mathbb{D}}c_{t}(x)\,d\mu(x)=\frac{-\partial_{t}\eta_{t}}{H_{t}}\,$ (23)
where
$\partial_{t}\eta_{t}=\int_{\mathbb{D}}\big(\gamma(\varphi_{t}(x))\big)\,\exp\!\Big(-\int_{0}^{t}\gamma\big(\varphi_{u}(x)\big)\,du\Big)\,y(x)\,d\mu(x),\qquad t\geq 0.$
Moreover, the population-weighted aggregate optimal portfolio in stocks is
$\pi_{t}^{\mu}\coloneqq\int_{\mathbb{D}}\pi_{t}(x)\,d\mu(x)=H_{t}^{-1}\eta_{t}\,\kappa_{t}-\varpi^{\mu}_{t},\qquad\varpi^{\mu}_{t}\coloneqq\int_{\mathbb{D}}\varpi_{t}(x)\,d\mu(x)$
where $\varpi^{\mu}_{t}$ is the aggregate hedging portfolio associated with $\varpi_{t}(x)$ that hedges the rate of endowment structure $Q$.
Next, we study the concept of equilibrium in our framework.
Definition 12.
Assume an economy $\mathcal{E}$ with underlying market $\mathfrak{M}$ and population structure $\boldsymbol{\mu}$ (with initial distribution $\mu$) as in Definition 4, and aggregate hedgeable (by state-tame portfolios) rate of consumption and rate of endowment structures $(\xi,c,Q)$ and current value of future endowments $L$ with non-arbitrage portfolio $\pi$. We say that $\mathcal{E}$ is an equilibrium if the following are satisfied for $\mu$:
Clearing of the money market:
$\xi^{\mu}_{t}-\pi^{\mu}_{t}=0$
for all $t\geq 0$, a.e.
Clearing of the commodity market:
$c^{\mu}_{t}=Q^{\mu}_{t}+D^{\mu}_{t}>0$
for all $0\leq t$ a.e., where $D^{\mu}_{t}=\delta^{\mu}_{t}P^{\mu}_{t}$.
Clearing of the stock market:
$\pi^{\mu}_{t}=P^{\mu}_{t}$
almost everywhere,
where $\xi^{\mu}_{t}$, $\pi^{\mu}_{t}$, $c^{\mu}_{t}$, $Q^{\mu}_{t}$, and $L^{\mu}_{t}$ are the aggregate processes associated with $\xi_{t}(x)$, $\pi_{t}(x)$, $c_{t}(x)$, $Q_{t}(x)$ and $L_{t}(x)$ respectively as in Definition 2.
Definition 13.
Assume an economy $\mathcal{E}$ with underlying market $\mathfrak{M}$ as defined by Theorem 2, a population structure $\boldsymbol{\mu}$ (with initial distribution $\mu$) and hedgeable (by state-tame portfolios) rate of consumption and rate of endowment structures $(\xi,c,Q)$, current value of future endowments $L$ and non-arbitrage portfolio $\pi$ defined by Theorem 2. We say that $\mathcal{E}$ is a short-horizon Duesenberry equilibrium if the equations and conditions defining the clearing of the money market, the clearing of the commodity market and the clearing of the stock market of Definition 12 hold.
Theorem 3.
Let $\varphi$ be a $\mathcal{C}(\mathbb{D}:\mathbb{D})$-valued Brownian flow of $C^{2}(\mathbb{D})$ diffeomorphisms, and let $\boldsymbol{\mu}$ be a population structure on $\mathbb{D}$, with $\mu$ initial population measure (see Definition 2). Assume a dynamic, consistent isoelastic preference structure for a population $U_{s,t}(x)$ as in Definition 10, with corresponding effective impatience rate $\gamma(x)$ of type $C^{2}(\mathbb{D})$ satisfying Assumption 2.
  1. (a)
    (Characterization.) Suppose $\mathcal{E}$ is a short-horizon Duesenberry equilibrium economy with underlying market $\mathfrak{M}$ with population structure $\boldsymbol{\mu}$ that satisfies the smooth market condition (Condition 1) with aggregate hedgeable (by state-tame portfolios) rate of consumption and rate of endowment structures $(\xi^{\mu},c^{\mu},Q^{\mu})$ and non-arbitrage portfolio $\pi^{\mu}$ defined by Theorem 2. Define the population-weighted effective aggregate process $\eta_{t}=\eta^{\mu}_{t}$
    $\eta_{t}\coloneqq\int_{\mathbb{D}}\exp\!\Big(-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du\Big)\,y(x)\,\,d\mu(x),$ (24)
    and the aggregate Duesenberry loading
    $\partial_{t}\eta_{t}=\int_{\mathbb{D}}\gamma(\varphi_{t}(x))\,\exp\!\Big(-\int_{0}^{t}\gamma\big(\varphi_{u}(x)\big)\,du\Big)\,y(x)\,d\mu(x),\qquad t\geq 0.$ (25)
    (since $\gamma$ is bounded above). Then, there exists a $C^{2}$ function $y(x)$ with
    $\int_{\mathbb{D}}y(x)\,\gamma(x)\,d\mu(x)=I^{\mu}_{0}=Q^{\mu}_{0}+D^{\mu}_{0}>0,$ (26)
    where $D^{\mu}_{t}=\delta^{\mu}_{t}P^{\mu}_{t}$, and $I_{t}^{\mu}=Q_{t}^{\mu}+D_{t}^{\mu}$ is the aggregate income process. Moreover, the state price process satisfies
    $H^{\mu}_{t}=\frac{-\partial_{t}\eta_{t}}{I^{\mu}_{t}},$ (27)
    and
    $P^{\mu}_{t}=\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}\,ds-\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\,\eta_{t}=\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\left(\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}\,ds-\eta_{t}\right),$ (28)
    where
    $Q^{\mu}_{t}=\int_{\mathbb{D}}Q^{\mu}_{t}(x)\,d\mu(x),$
    and $\Psi^{\mu}_{t}$ is defined as
    $\Psi^{\mu}_{t}=-E\!\left[\frac{Q^{\mu}_{t}}{I^{\mu}_{t}}\,\partial_{t}\eta_{t}\right]=\int_{\mathbb{D}}\Psi^{\mu}_{t}(x)\,d\mu(x),$ (29)
    with
    $\Psi^{\mu}_{t}(x)=-E\!\left[\frac{Q^{\mu}_{t}(x)}{I^{\mu}_{t}}\,\partial_{t}\eta_{t}\right]=E\!\left[H^{\mu}_{t}\,Q_{t}^{\mu}(x)\right].$
  2. (b)
    (Existence.) Assume non-negative (semimartingale) random fields $Q_{t}^{\mu}(x)$ and $I_{t}^{\mu}(x)$ of type $C^{2}(\mathbb{D})$ consistent with the population structure $\boldsymbol{\mu}$, satisfying (either) Assumption 3 or Assumption 4. Assume $0\leq Q_{t}^{\mu}(x)<I_{t}^{\mu}(x)$ with $I_{t}^{\mu}(x)>0$ for all $x\in\mathbb{D}$, and $t\geq 0$. Moreover, assume that there exists a $C^{2}$ function $y(x)>0$ for all $x$ with
    $0<\int_{\mathbb{D}}y(x)\gamma(x)\,d\mu(x)=\int_{\mathbb{D}}I_{0}^{\mu}(x)\,d\mu(x)<\infty,$ (30)
    Furthermore, we define $\Psi^{\mu}_{t}(x)$ and $\Psi^{\mu}_{t}$ by the following equations:
    $\Psi^{\mu}_{t}(x)=-E\!\left[\frac{Q^{\mu}_{t}(x)}{I_{t}^{\mu}}\,\partial_{t}\eta_{t}\right],\qquad\Psi^{\mu}_{t}=\int_{\mathbb{D}}\Psi^{\mu}_{t}(x)\,d\mu(x)$ (31)
    where
    $\int_{0}^{\infty}\int_{\mathbb{D}}\Psi^{\mu}_{t}(x)\,d\mu(x)\,dt<\infty\qquad\text{and}\qquad\int_{0}^{\infty}\Psi^{\mu}_{t}(x)\,dt<\infty,\qquad\forall x\in\mathbb{D}.$ (32)
    We define the market $\mathfrak{M}=(P^{\mu},\boldsymbol{\mu},b^{\mu},\sigma^{\mu},\delta^{\mu},\vartheta^{\mu},r^{\mu})$ by the following equations:
    $H^{\mu}_{t}=\frac{-\partial_{t}\eta_{t}}{I_{t}^{\mu}},$ (33)
    and
    $P^{\mu}_{t}=\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\left(\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}\,ds-\eta_{t}\right),\qquad\delta_{t}^{\mu}P^{\mu}_{t}=D_{t}^{\mu}=I_{t}^{\mu}-Q_{t}^{\mu},$ (34)
    where $I^{\mu}_{t}$ and $Q^{\mu}_{t}$ are the aggregate processes of $I_{t}^{\mu}(x)$, and $Q_{t}^{\mu}(x)$ respectively, and it is assumed that there exists a continuous semimartingale process $\kappa^{\mu}_{t}$, consistent with the population structure $\boldsymbol{\mu}$, with
    $\kappa^{\mu}_{t}\sigma^{\mu}_{t}=\vartheta^{\mu}_{t}.$ (35)
    Then, there exists a short-horizon Duesenberry equilibrium $\mathcal{E}$ with underlying market $\mathfrak{M}$ with population structure $\boldsymbol{\mu}$ and a dynamic, consistent isoelastic preference structure for a population with effective impatience rate $\gamma(x)$ with type space $\mathbb{D}$.
  3. (c)
    In either case, the underlying market is free of state-tame arbitrage.
Proof of Theorem 3.
We divide the argument into two parts. First, we prove the characterization: starting from a short-horizon Duesenberry equilibrium, we obtain the identities (24)–(26) and the pricing formulas (27)–(28). Then, we prove the existence direction by reversing the construction. Finally, we prove non-arbitrage of the market in both cases.
Characterization. Assume a short-horizon Duesenberry equilibrium $\mathcal{E}$ with underlying market $\mathfrak{M}$ that satisfies the Smooth Market Condition (Condition 1). Under the hypotheses of Theorem 2, for each agent of (initial) type $x\in\mathbb{D}$, the locally optimal wealth, consumption, and portfolio processes are given by $(\xi^{\mu},c^{\mu},Q^{\mu})$ and $\pi^{\mu}$ as follows:
$\xi^{\mu}_{t}(x)=L^{\mu}_{t}(x)+e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}(H^{\mu}_{t})^{-1}y(x)$ (36)
where the initial net wealth $y(x)=\xi^{\mu}_{0}(x)-L^{\mu}_{0}(x)$ is a function of class $C^{2}$ and
$c^{\mu}_{t}(x)=\gamma(\varphi_{t}(x))e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}(H^{\mu}_{t})^{-1}y(x)$ (37)
and a state-tame (no-arbitrage) portfolio in the sense of Theorem 2 and Remark 1 given by:
$\pi^{\mu}_{t}(x)=e^{-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du}\,(H^{\mu}_{t})^{-1}y(x)\,\kappa^{\mu}_{t}\;-\;\varpi^{\mu}_{t}(x).$ (38)
Now, we aggregate over the (random) population at time $t$ the second term of the right-hand side of Equation (36). This yields Equation (24) by Corollary 2. Similarly, integrating equations (36) and (37) over the (time-$t$) distribution $\mu_{t}$ gives the aggregate processes
$\xi_{t}^{\mu}=L_{t}^{\mu}\;+\;\big(H_{t}^{\mu}\big)^{-1}\eta_{t}$ (39)
and
$\partial_{t}\eta_{t}=\int_{\mathbb{D}}\gamma\big(\varphi_{t}(x)\big)\,\exp\!\Big(-\int_{0}^{t}\gamma\big(\varphi_{u}(x)\big)\,du\Big)\,y(x)\,d\mu(x).$ (40)
We notice that $t\mapsto\exp(-\int_{0}^{t}\gamma(\varphi_{u}(x))du)$ is non-increasing for each $x$, hence, so is $\eta_{t}$. Also, by Corollary 4 $\eta_{t}$, and $\partial_{t}\eta_{t}$ are consistent semimartingales with respect to the population structure $\boldsymbol{\mu}$. Moreover, by the dominated convergence theorem $\eta_{t}\to 0$ a.s.
Aggregating (37) against the population structure gives
$c_{t}^{\mu}=-\big(H^{\mu}_{t}\big)^{-1}\,\partial_{t}\eta_{t}.$ (41)
By the clearing of the commodity market $c_{t}^{\mu}=I_{t}^{\mu}>0$ for all $t\geq 0$ (a.e.), we obtain
$H^{\mu}_{t}=\frac{-\partial_{t}\eta_{t}}{I^{\mu}_{t}},$ (42)
which is Equation (27), where $Q_{t}^{\mu}$ is the aggregate (hedgeable) endowment rate and $D^{\mu}_{t}=\delta^{\mu}_{t}P^{\mu}_{t}$ is the total dividends of the aggregate stock. Using the clearing of the commodity market at $t=0$ and using (40) we obtain Equation (26).
We notice that by Equation (42), the dominated convergence theorem, Assumption 2, and Equation (26), imply that
$0\leq\int_{0}^{\infty}H^{\mu}_{t}Q_{t}^{\mu}\,dt=\int_{0}^{\infty}\frac{Q^{\mu}_{t}}{I^{\mu}_{t}}(-\partial_{t}\eta_{t})\,dt\leq\int_{0}^{\infty}(-\partial_{t}\eta_{t})\,dt=\eta_{0}<\infty$ (43)
By Equation (43), since $(L^{\boldsymbol{\mu}},Q^{\boldsymbol{\mu}})$ is a hedgeable rate of endowment (with non-arbitrage portfolio $\pi^{\boldsymbol{\mu}}_{Q}$), using Fubini’s theorem and the cocycle property, it follows that
$L_{t}^{\mu}=\int_{\mathbb{D}}L_{t}(x)\,d\mu(x)=-\,(H^{\mu}_{t})^{-1}\,E\!\left[\int_{t}^{\infty}H^{\mu}_{u}\,Q_{u}^{\mu}\,du\;\Big|\;\mathcal{F}_{t}\right]=-\,(H^{\mu}_{t})^{-1}\,\int_{t}^{\infty}\Psi^{\mu_{t}}_{u}\,du$ (44)
Equation (28) is a direct consequence of Corollary 2, and Equation (44).
This completes the derivation of (28) from the market-clearing conditions together with (27).
On the other hand, by Itô’s Lemma, Equation (27), Equation (28), and since $\eta_{t}\to 0$ it follows that
$E\Big[\int_{0}^{\infty}\big(1-\frac{Q^{\mu}_{u}}{I_{u}^{\mu}}\big)(-\partial_{u}\eta_{u})\,du\mid\mathcal{F}_{t}\Big]=H_{t}^{\mu}P_{t}^{\mu}+\int_{0}^{t}H_{u}^{\mu}D_{u}^{\mu}\,du=\\ P_{0}^{\mu}+\int_{0}^{t}H_{u}^{\mu}P_{u}^{\mu}\Big(b^{\mu}_{u}+\delta_{u}^{\mu}-r^{\mu}_{u}-(\sigma_{u}^{\mu})^{\intercal}\vartheta_{u}^{\mu}\Big)\,du+\int_{0}^{t}H^{\mu}_{u}P_{u}^{\mu}\left((\sigma^{\mu}_{u})^{\intercal}-(\vartheta^{\mu}_{u})^{\intercal}\right)dW(u)$
It follows that $b^{\mu}_{u}+\delta_{u}^{\mu}-r^{\mu}_{u}-(\sigma_{u}^{\mu})^{\intercal}\vartheta_{u}^{\mu}=0$, proving that there are no state-tame arbitrage opportunities.
Existence. Assume non-negative random fields $Q^{\mu}(x)$ and $I_{t}^{\mu}(x)$ of type $C^{2}(\mathbb{D})$, with the property that there exist a $C^{2}$ function $y(x)>0$ for all $x$ satisfying Equation (26). Define $\Psi^{\mu}_{t}(x)$, and $\Psi^{\mu}_{t}$ by Equation (31), assuming Equation (32). Define, the effective aggregate process $\eta_{t}>0$ by Equation (24), and define $H_{t}^{\mu}$ by Equation (33). By (26) and (24),
$\int_{\mathbb{D}}\gamma(x)\,y(x)\,d\mu(x)=I_{0}^{\mu},$
hence $H^{\mu}_{0}=1$ by Equation (27). Since $\gamma(x)>0$, $y(x)>0$ and $I^{\mu}_{t}(x)>0$ for all $x$, $H^{\mu}_{t}>0$.
Let $(r^{\mu},\vartheta^{\mu})$ be the coefficients obtained from the Itô decompositions of $H^{\mu}$. Also,
$\int_{t}^{\infty}\Psi^{\mu}_{s}\,ds=E\big[\int_{0}^{\infty}\int_{\mathbb{D}}\frac{Q^{\mu}_{u}(x)}{I^{\mu}_{u}}(-\partial_{u}\eta_{u})\,d\mu(x)\,du\Big]<\\ E\Big[\int_{0}^{\infty}\Big(-\partial_{u}\eta_{u}\Big)\,du\Big]=E[\eta_{0}]<\infty$
Moreover,
$\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}\,ds-\eta_{t}=E\Big[\int_{t}^{\infty}\frac{Q_{u}^{\mu}}{I^{\mu}_{u}}(-\partial_{u}\eta_{u})\,du-\int_{t}^{\infty}(-\partial_{u}\eta_{u})\,du\mid\mathcal{F}_{t}\Big]<0$ (45)
It follows from Equation (34), and Equation (45):
$P^{\mu}_{t}=\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\left(\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}\,ds-\eta_{t}\right)>0.$
Define $(b^{\mu},\sigma^{\mu})$ the coefficients obtained satisfying Assumption 3 or 4 according to the underlying hypothesis of the theorem, and define $\delta_{t}^{\mu}\coloneqq D^{\mu}_{t}/P^{\mu}_{t}$ as the aggregate dividend yield, where $D_{t}^{\mu}=I_{t}^{\mu}-Q_{t}^{\mu}>0$ is the aggregate dividend income. Assume that (35) holds with a consistent (relative to the population structure $\boldsymbol{\mu}$) process $\kappa^{\mu}_{t}$ such that $\kappa^{\mu}_{t}\sigma^{\mu}_{t}=\vartheta^{\mu}_{t}$. In particular the market defined by $P_{t}^{\mu}$ and $H_{t}^{\mu}$ above satisfies the smooth market condition, and therefore Theorem 2 and Corollary 2 apply. Consider the consumer problem in the market $\mathfrak{M}=(P^{\mu},\boldsymbol{\mu},b^{\mu},\sigma^{\mu},\delta^{\mu},\vartheta^{\mu},r^{\mu})$ as above, with the given consistent isoelastic preference structure. Define $\big(\xi^{\mu}(x),c^{\mu}(x),\pi^{\mu}(x)\big)$ the formulas of Theorem 2 satisfying the explicit formulas (36)–(37) (and the corresponding portfolio representation), with initial net wealth $y^{\mu}(x)=y(x)$, and initial financial wealth $\xi^{\mu}_{0}(x)=y(x)+L_{0}^{\mu}(x)$, and labor income $Q_{t}^{\mu}(x)$, so
$L^{\mu}_{t}(x)=\frac{I^{\mu}_{t}}{\partial_{t}\eta_{t}}\int_{t}^{\infty}\Psi^{\mu_{t}}_{s}(x)\,ds=(H_{t}^{\mu})^{-1}E\!\left[\int_{t}^{\infty}H^{\mu}_{u}\,Q_{u}^{\mu}(x)\,du\;\Big|\;\mathcal{F}_{t}\right],$
is the current value of future labor at time $t$ in the constructed market. It follows by Corollary 2, that $\xi_{t}^{\mu}=L_{t}^{\mu}+(H_{t}^{\mu})^{-1}\eta_{t}$, and therefore by Equation (34), $P_{t}^{\mu}=\xi_{t}^{\mu}$. Also, by Itô’s rule, and aggregation, it follows from Equation (23)
$H_{t}^{\mu}L_{t}^{\mu}-\int_{0}^{t}H^{\mu}_{u}Q_{u}^{\mu}\,du=\\ -E\Big[\int_{0}^{\infty}H^{\mu}_{u}Q^{\mu}_{u}\,du\Big]+\int_{0}^{t}H_{u}^{\mu}\Big((\pi_{Q}^{\mu})_{u}(\sigma_{u}^{\mu})^{\intercal}-L_{u}^{\mu}(\vartheta^{\mu}_{u})^{\intercal}\Big)dW(u),$ (46)
where $\pi_{Q}^{\mu}$ is the aggregate portfolio (that hedges $L^{\mu}_{t}$) implied by the representation as local martingale of the term on the left of the last equation.
We observe that
$H_{t}^{\mu}P_{t}^{\mu}+\int_{0}^{t}H_{u}^{\mu}D_{u}^{\mu}\,du=\\ E\Big[\Big(\int_{t}^{\infty}(-\partial_{u}\eta_{u})\,du-\int_{t}^{\infty}\frac{Q_{u}^{\mu}}{I^{\mu}_{u}}(-\partial_{u}\eta_{u})\,du\Big)\mid\mathcal{F}_{t}\Big]+\int_{0}^{t}\Big(I_{u}^{\mu}-Q_{u}^{\mu}\Big)\frac{(-\partial_{u}\eta_{u})}{I^{\mu}_{u}}\,du=\\ E\Big[\Big(\int_{0}^{\infty}(-\partial_{u}\eta_{u})\,du-\int_{0}^{\infty}\frac{Q_{u}^{\mu}}{I^{\mu}_{u}}(-\partial_{u}\eta_{u})\,du\Big)\mid\mathcal{F}_{t}\Big]$
is a martingale. On the other hand, Itô’s Lemma implies
$H_{t}^{\mu}P_{t}^{\mu}+\int_{0}^{t}H_{u}^{\mu}D_{u}^{\mu}\,du=P_{0}^{\mu}+\int_{0}^{t}H_{u}^{\mu}P_{u}^{\mu}\Big(b^{\mu}_{u}+\delta_{u}^{\mu}-r^{\mu}_{u}-(\sigma_{u}^{\mu})^{\intercal}\vartheta_{u}^{\mu}\Big)\,du+\\ \int_{0}^{t}H^{\mu}_{u}P_{u}^{\mu}\left((\sigma^{\mu}_{u})^{\intercal}-(\vartheta^{\mu}_{u})^{\intercal}\right)dW(u)$ (47)
Since the left-hand side of Equation (47) is a martingale, it follows that the bounded variation part is 0, implying that
$b_{u}^{\mu}+\delta_{u}^{\mu}-r_{u}^{\mu}=(\sigma_{u}^{\mu})^{\intercal}\vartheta_{u}^{\mu}$
and then the market, as defined above, is free of state-tame arbitrage opportunities. From Theorem 2, and Corollary 2 and Equation (33), it follows that
$c^{\mu}_{t}=\frac{-\partial_{t}\eta_{t}}{H^{\mu}_{t}}=I_{t}^{\mu}$ (48)
Moreover from Theorem 2, and Corollary 2, and Equation (48) it follows that
$H^{\mu}_{t}\xi_{t}^{\mu}+\int_{0}^{t}H_{u}^{\mu}(c_{u}^{\mu}-Q_{u}^{\mu})\,du=\xi_{0}+\int_{0}^{t}H_{u}^{\mu}\Big(\pi_{u}^{\mu}(\sigma_{u}^{\mu})^{\intercal}-\xi_{u}^{\mu}(\vartheta_{u}^{\mu})^{\intercal}\Big)\,dW(u)$ (49)
Taking differences of equations (49) and (47), and since $\xi_{t}^{\mu}=P_{t}^{\mu}$, and since $D^{\mu}_{t}=I_{t}^{\mu}-Q^{\mu}_{t}=c^{\mu}_{u}-Q^{\mu}_{u}$ by equation  (48) it follows that $P^{\mu}_{u}=\pi_{u}^{\mu}$, proving the clearing of the money market, and the clearing of the stock market. This completes the existence part and therefore the proof. ∎
Corollary 3.
Assume the conditions of Theorem 3. Let
$\eta_{t}:=\int_{\mathbb{D}}\exp\!\left(-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du\right)\,y(x)\,d\mu(x),$ (50)
be the population-weighted effective aggregate process. Then the state price process and the consumption–wealth ratio satisfy
$H_{t}=\frac{-\partial_{t}\eta_{t}}{c_{t}},\qquad\frac{c_{t}}{P^{W}_{t}}=\frac{-\partial_{t}\eta_{t}}{\eta_{t}}=-\partial_{t}\log(\eta_{t}),$ (51)
where $c_{t}=c_{t}^{\mu}$ is aggregate consumption and
$P_{t}^{L}:=(H_{t})^{-1}E\!\left[\int_{t}^{\infty}H_{u}\,Q_{u}^{\mu}\,du\;\middle|\;\mathcal{F}_{t}\right]=-\,L_{t}^{\mu},\qquad P_{t}^{W}:=P_{t}+P_{t}^{L}=(H_{t})^{-1}\eta_{t}$
is the total wealth (physical wealth plus labor wealth) owned by the society at time $t$. Moreover, since $\eta$ is a.s. absolutely continuous in $t$, the volatility $\sigma_{t}^{W}$ of $P_{t}^{W}$ is equal to the market price of risk $\vartheta_{t}^{\mu}$.
Remark 7.
The right-hand side of Equation (51) implies that the consumption–wealth ratio is the instantaneous proportional decay rate of $\eta_{t}$, where $\eta_{t}$ represents the impatience-adjusted social wealth: it is a cross–sectional average of initial wealth $y(x)$, weighted by the cumulative impatience factor along each trajectory through the effective impatience field $\gamma(\varphi_{u}(x))$. A consequence of the latter is that the only factor that accounts for changes in the consumption–wealth ratio is derived from changes in the preference structure of the population, and not from factors such as productivity, population growth, or inflation, at least directly. See Lustig et al. [49] for empirical analysis of the wealth–consumption ratio, and Lettau and Ludvigson [39] for the closely related cay variable, which has been shown to forecast stock returns and whose fluctuations, in the present framework, are driven entirely by shifts in the cross-sectional impatience structure $\eta_{t}$.
Volatility decomposition. If $\sigma^{c}_{t}$ denotes the volatility of aggregate consumption (relative diffusion) and
$\sigma^{-\partial\eta}_{t}=\frac{1}{-\partial_{t}\eta_{t}}\int_{\mathbb{D}}\big(\nabla\gamma(\varphi_{t}(x))\varrho(\varphi_{t}(x))\big)\,\exp\!\Big(-\int_{0}^{t}\gamma\big(\varphi_{u}(x)\big)\,du\Big)\,y(x)\,d\mu(x)$
is the volatility (relative diffusion) of $\partial_{t}\eta_{t}$, then the volatility $\sigma_{t}^{W}$ of total wealth $P^{W}_{t}$ (relative diffusion) satisfies
$\sigma_{t}^{W}=\vartheta_{t}=\sigma_{t}^{c}-\sigma_{t}^{-\partial\eta},$
where $\vartheta_{t}$ is the market price of the risk process. Hence, asset-price risk premia and market volatility (which are governed by the diffusion of the state-price process) need not be tied to consumption volatility: even if aggregate consumption is smooth (small $\|\sigma_{t}^{c}\|$), aggregate wealth (or the state-price) can be highly volatile whenever impatience risk generates large fluctuations in $\partial_{t}\eta_{t}$. This provides a direct valuation channel through which time-varying effective impatience can reconcile low consumption volatility with volatile markets. We note that this mechanism is distinct from the long-run risk channel of Bansal and Yaron [6], where persistent shocks to consumption growth drive asset-price volatility, and from the external habit mechanism of Campbell and Cochrane [11], where a time-varying surplus ratio amplifies risk premia. In our framework, the additional source of market volatility arises from heterogeneous and time-varying impatience across the population, encoded in the term $\sigma_{t}^{-\partial\eta}$.
Remark 8.
Equity premium decomposition. The volatility decomposition $\vartheta_{t}=\sigma_{t}^{c}-\sigma_{t}^{-\partial\eta}$ of Remark 7, together with the no-arbitrage relation (8), yields an explicit decomposition of the equity premium of any traded asset (or portfolio) with price process $P^{\Sigma}_{t}$, rate of return $b_{t}^{\Sigma}$, dividend yield $\delta_{t}^{\Sigma}$, and volatility vector $\sigma_{t}^{\Sigma}$:
$b_{t}^{\Sigma}+\delta_{t}^{\Sigma}-r_{t}\;=\;\underbrace{(\sigma_{t}^{c})^{\intercal}\,\sigma_{t}^{\Sigma}\,}_{\displaystyle\text{consumption risk premium}}\;-\;\underbrace{(\sigma_{t}^{-\partial\eta})^{\intercal}\,\sigma_{t}^{\Sigma}\,}_{\displaystyle\text{impatience risk premium}}$ (52)
The first term is the classical Consumption-CAPM premium (see Breeden [9]): the covariance of aggregate consumption growth with asset returns. The second term is the impatience risk premium: the covariance of the growth of the Duesenberry loading $\partial_{t}\eta_{t}$ with asset returns.
From market volatility to aggregate wealth volatility. A convenient way to read (52) is that equity premia are governed by an inner product of volatilities. In fact, as a consequence of Equation (51), the market price of risk $\vartheta_{t}=\sigma^{W}_{t}$, where $\sigma^{W}_{t}$ is the volatility of total wealth (physical wealth plus labor wealth) $P_{t}^{W}$, and therefore Equation (52) can be rewritten as
$b_{t}^{\Sigma}+\delta_{t}^{\Sigma}-r_{t}=(\sigma_{t}^{W})^{\intercal}\sigma_{t}^{\Sigma},$
where $\sigma_{t}^{\Sigma}$ denotes the market (equity) volatility. In particular, the proxy $\sigma^{W}\approx\sigma^{\Sigma}$ would predict an equity premium of order $(\sigma^{\Sigma})^{2}$. Table 1 collects long-run U.S. estimates from three independent sources and compares the observed equity premium with the value $\widehat{\mathrm{EP}}\coloneqq(\sigma^{\Sigma})^{2}$ predicted by this proxy.
Table 1. Observed versus predicted equity premium under the proxy $\sigma^{W}\!\approx\sigma^{\Sigma}$: long-run U.S. estimates.
Source Period $\sigma^{\Sigma}$ $\widehat{\mathrm{EP}}=(\sigma^{\Sigma})^{2}$ $\mathrm{EP}$ $\vartheta$
Mehra [51] 1889–2000 $0.186$ $3.5\%$ $6.9\%$ $0.37$
Jorda et al. [31] 1870–2015 $0.192$ $3.7\%$ $6.94\%$ $0.36$
Jorda et al. [31] 1963–2015 $0.163$ $2.7\%$ $5.69\%$ $0.35$
Campbell [10] 1891–1995 $0.186$ $3.5\%$ $4.74\%$ $0.25$
Campbell [10] 1947–1996 $0.155$ $2.4\%$ $6.78\%$ $0.44$
Campbell [10] 1970–1996 $0.174$ $3.0\%$ $4.54\%$ $0.26$
Notes. $\sigma^{\Sigma}$ is the annualized standard deviation of real equity returns. $\widehat{\mathrm{EP}}=(\sigma^{\Sigma})^{2}$ is the equity premium predicted by the proxy $\sigma^{W}\approx\sigma^{\Sigma}$. $\mathrm{EP}$ is the observed mean excess return of equities over bills, annualized. $\vartheta\coloneqq\mathrm{EP}/\sigma^{\Sigma}$ is the implied market price of risk, which under the equilibrium coincides with $\sigma^{W}$. In every sample $\mathrm{EP}>\widehat{\mathrm{EP}}$, equivalently $\vartheta>\sigma^{\Sigma}$. Data for Mehra [51]: Table 1 and Sharpe-ratio calculation on p. 60 therein (S&P 500, 1889–2000). Jorda et al. [31]: Tables 3–5 (U.S. equity excess returns over bills; full and post-WW2 balanced panels). Campbell [10]: Table 2 therein (log returns on a broad U.S. stock index over short-term government debt).
Our framework does not require aggregate wealth to be “as smooth as equity”. The equilibrium identifies $\vartheta_{t}$ with a volatility-type object that naturally admits an interpretation as an aggregate wealth volatility. As Table 1 confirms, empirical estimates of $\vartheta$ range from 0.250.250.370.37 in long-run U.S. data (Mehra 51; Jorda et al. 31; Campbell 10) to roughly $0.44$ in the post-war quarterly sample of Campbell 10, implying an unobserved volatility of total wealth of the same order; the “puzzle” is shifted from consumption volatility to the volatility of (partly unobserved) aggregate wealth.
Connection to credit conditions and a forthcoming extension. Table 1 reveals that $\vartheta=\sigma^{W}$ consistently exceeds $\sigma^{\Sigma}$ in every sample, implying that the volatility of aggregate wealth is substantially larger than market volatility alone. A natural explanation arises from the composition of total wealth $P_{t}^{W}=P_{t}+P_{t}^{L}$: the present value of future labor income $P_{t}^{L}$ is discounted, in the model, at the market interest rate $r_{t}$. Moreover, the effective rates at which households actually discount future consumption streams—consumer credit, credit-card, and unsecured borrowing rates—are both significantly higher and considerably more volatile than the risk-free rate. Because $P_{t}^{L}$ typically accounts for two-thirds or more of aggregate wealth (Lustig et al. [49]), even moderate fluctuations in these consumer-relevant discount rates translate into large swings in $P_{t}^{W}$, generating an aggregate wealth volatility $\sigma^{W}$ well in excess of equity volatility $\sigma^{\Sigma}$.
The above remark suggests an extension—Duesenberry equilibria with endogenous borrowing premia [48]—in which the relevant discounting of labor cash-flows is performed at a household borrowing rate (or, more generally, a state-dependent effective rate generated by a hazard/intensity mechanism) rather than at $r_{t}$. However, as shown in our related analysis (see Londoño [47]), introducing a state-dependent adjustment to the effective borrowing rate—interpretable as a hazard intensity $\lambda(\varphi_{t}(x))$ reflecting default and liquidity risks—does not modify the aggregate pricing kernel $H_{t}$ nor the optimal consumption path (expressed relative to the relevant present value). It only re-evaluates the collateralizable present value of labor income entering the borrowing constraint. Hence, large fluctuations in total wealth arise from state-dependent discounting of future labor flows, without altering aggregate risk pricing or consumption dynamics. We develop this extension in a forthcoming work (see Londoño [48]).
Interest rate decomposition. A parallel decomposition holds for the short rate. Applying Itô’s formula to $H_{t}=(-\partial_{t}\eta_{t})/c_{t}$ and writing $dc_{t}/c_{t}=\mu_{t}^{c}\,dt+(\sigma_{t}^{c})^{\intercal}dW_{t}$ and $d(-\partial_{t}\eta_{t})/(-\partial_{t}\eta_{t})=\mu_{t}^{-\partial\eta}\,dt+(\sigma_{t}^{-\partial\eta})^{\intercal}dW_{t}$ for the relative drifts and diffusions of aggregate consumption and the Duesenberry loading, respectively, one obtains
$r_{t}\;=\;\bigl(\mu_{t}^{c}-\mu_{t}^{-\partial\eta}\bigr)\;-\;(\sigma_{t}^{c})^{\intercal}\,\vartheta_{t}\,.$ (53)
The first term captures the net drift of aggregate consumption over the Duesenberry loading; the second is a precautionary-savings correction that depends on the covariance of consumption growth with the market price of risk. In the constant-impatience case, $\mu_{t}^{-\partial\eta}=-\gamma$ and (53) reduces to $r_{t}=\mu_{t}^{c}+\gamma-\|\sigma_{t}^{c}\|^{2}$, the standard expression.
[…]
The source of the puzzle is transparent in (53): when the population is homogeneous, $\vartheta_{t}=\sigma_{t}^{c}$ and the precautionary-savings correction reduces to $(\sigma_{t}^{c})^{\intercal}\vartheta_{t}=(\sigma_{t}^{c})^{\intercal}\sigma_{t}^{c}\approx 0.02\%$, a negligible reduction. In the heterogeneous-impatience economy, however, $\vartheta_{t}=\sigma_{t}^{c}-\sigma_{t}^{-\partial\eta}$ is the full market price of risk (Remark 7), so the correction becomes
$(\sigma_{t}^{c})^{\intercal}\vartheta_{t}\;=\;\|\sigma_{t}^{c}\|^{2}\;-\;(\sigma_{t}^{c})^{\intercal}\sigma_{t}^{-\partial\eta},$
which is substantially amplified whenever $(\sigma_{t}^{c})^{\intercal}\sigma_{t}^{-\partial\eta}<0$, that is, when consumption growth and the Duesenberry loading co-vary negatively. Using the empirical consumption risk premium estimate $(\sigma^{c})^{\intercal}\vartheta\approx 2.38\%$ of Lustig et al. [49] and retaining $\mu^{-\partial\eta}\approx-\gamma$ as the leading-order drift, Equation (53) gives
$r\;\approx\;\underbrace{(\mu^{c}+\gamma)}_{\approx\,6.75\%}\;-\;\underbrace{(\sigma^{c})^{\intercal}\vartheta}_{\approx\,2.38\%}\;\approx\;4.37\%,$
[…]
Remark 9.
Comparison with competing approaches to the equity premium puzzle. The decomposition (52) and the calibration exercise above relies exclusively on market observables: the Sharpe ratio $\|\vartheta\|$, aggregate consumption volatility $\|\sigma^{c}\|$, and the consumption risk premium $(\sigma^{c})^{\intercal}\vartheta$, and the equity volatility $\sigma^{\Sigma}$. No preference parameter needs to be calibrated beyond the requirement that the cross-sectional impatience distribution be non-degenerate. This stands in sharp contrast to the leading approaches in the literature, each of which introduces at least one latent variable or a poorly identified parameter to generate the observed equity premium.
(i) Representative-agent CRRA models (Jorda et al. [31], Campbell [10]) require a coefficient of relative risk aversion for equities of at least 37374848 (on the post-WW2 U.S. sample, depending on sample and data source), far above the range considered economically plausible; our decomposition nests this case with $\sigma^{-\partial\eta}=0$, so the failure of the CRRA model is precisely the absence of the impatience risk premium. (ii) The external habit model of Campbell and Cochrane [11] matches aggregate moments but relies on a reverse-engineered sensitivity function $\lambda(S_{t})$, entailing anomalous welfare properties: Ljungqvist and Uhlig [40] demonstrate that destroying part of the endowment can raise welfare under the Campbell–Cochrane specification. (iii) The long-run risk model (Bansal and Yaron [6]; Epstein and Zin [19]) rests on a latent persistent component in consumption growth whose implied predictability is not found in the data (Beeler and Campbell [7]), a disputed elasticity of intertemporal substitution (Hall [28]; Yogo [66]), and an implausibly strong preference for early resolution of uncertainty (Epstein et al. [20]). (iv) The incomplete-markets heterogeneity of Constantinides and Duffie [13] is an existence result whose quantitative assessment requires countercyclically heteroscedastic idiosyncratic shocks of debated magnitude (Kocherlakota [35]). Brav et al. [8] show that the channel can account for the equity premium when household-level consumption data are used, but this requires micro-level consumption panels that are unavailable in most countries and do not settle the question at the aggregate level.
In each case, the additional state variable is either latent, governed by poorly identified parameters, or tied to preference specifications with implausible side effects. The Duesenberry equilibrium requires none of these ingredients: $\sigma^{-\partial\eta}$ is not a free parameter, but an identity implied by the equilibrium conditions (51), fully determined once $\vartheta$ and $\sigma^{c}$ are measured using market data.
Remark 10.
Inflation and the nominal short rate. The left-hand side of Equation (51), combined with the dynamics
$dH_{t}=-r_{t}H_{t}\,dt-\vartheta_{t}^{\intercal}H_{t}\,dW_{t},$
implies that if inflation does not alter the impatience parameter of the population, then inflation is transmitted linearly into the nominal short rate in the long run (since it only enters $H_{t}$ through aggregate consumption $c_{t}^{\mu}=I_{t}^{\mu}$). However, the transmission of inflation to the nominal short rate need not be purely linear, since a pure price-level effect rescales aggregate nominal consumption, and inflation might generate short-run distributional wedges because nominal labor income adjusts sluggishly (relative to prices and aggregate consumption). In the present framework, this sluggishness matters because (51) implies that inflation can affect not only the denominator $c_{t}$ but also the numerator $\partial_{t}\eta_{t}$ through changes in effective impatience and wealth/income weights. Hence, deviations from a one-for-one mapping can be interpreted as equilibrium adjustments in the impatience-adjusted wealth component induced by inflation-driven income misalignment. See Fisher [22], Mishkin [53] for the Fisher effect and its empirical evidence. See also Auclert [5], Doepke and Schneider [14] for literature studying the effect of inflation on the distribution of wealth (see also Erosa and Ventura [21] for an equilibrium analysis of inflation as a regressive consumption tax), and see Mian et al. [52], Rannenberg [59], Rachel and Summers [57], among others, for studies on the effect of wealth distribution on interest rates.
Population growth. As a consequence of the left-hand side of Equation (51), population growth affects aggregates through the population structure (hence through $c_{t}=I_{t}^{\mu}$) in a weighted manner, reflecting income weights. In particular, assuming an absolutely continuous population growth structure as in this paper, its aggregation weight by wealth contributes linearly to the short-term interest rate and does not alter the volatility of aggregate consumption or the market volatility. See Carvalho et al. [12], Gagnon et al. [26], Aksoy et al. [2] for references relating demographic changes to changes in the short interest rate, and Eggertsson et al. [18] for a quantitative equilibrium analysis of how population aging depresses the natural rate.
Multiplicative invariance. Any purely multiplicative scaling of the aggregate income process (from inflation or population size) rescales $c_{t}$ and $P_{t}^{W}$ in the same proportion, while leaving $c_{t}/P_{t}^{W}=-\partial_{t}\log(\eta_{t})$ unchanged.
5. Examples
The purpose of this section is to illustrate two complementary (and essentially equivalent) ways to parametrize tractable short-horizon Duesenberry equilibria under smooth market conditions.
Example 2 is most useful when the modeler wishes to treat the valuation of human wealth (the present value of labor) as the endogenous primitive, whereas wages are a derived quantity.
Example 3 reverses the construction and is tailored to situations where the labor income process is the endogenous primitive.
We assume the following for this section: Assume $\mathbb{D}\subset\mathbb{R}^{d}$ and let $(\varphi_{t})_{t\geq 0}$ be a common Brownian flow of $C^{2}$–diffeomorphisms generated by
$d\varphi_{t}(x)={\rho}(\varphi_{t}(x))\,dt+\varrho(\varphi_{t}(x))\,dW_{t}\qquad\varphi_{0}(x)=x,$
where $W$ is an $n$-dimensional Brownian motion and the coefficients ensure a non-explosive strong solution and Kunita–type regularity. Assume a population structure $\boldsymbol{\mu}$ with population growth rate $h$, and assume an isoelastic consistent preference structure for a population $U$ with effective impatience field $\gamma(\varphi_{t}(x)(\omega))$ of $U_{t}(\omega)(x)$, where $\gamma(\cdot)$ is a function of type $C^{2}(\mathbb{D})$, satisfying Assumption 2. Assume a positive income function $I\in C^{2}(\mathbb{D})$, and an initial total wealth $y$ of type $C^{2}$ with $y(x)>0$ for all $x$ satisfying Equation (30). Define the type-dependent income and aggregate income as
$I_{t}(x)\coloneqq\Lambda_{t}(x)I(\varphi_{t}(x)),\qquad I_{t}^{\mu}=\int_{\mathbb{D}}I_{t}(x)\,d\mu(x)$
where the population weights $\Lambda_{s,t}(x)$ as in Definition 2 and the (finite-variation) kernel are given by
$\Lambda_{t}(x)=\Lambda_{0,t}(x)=\exp\!\Big(\int_{0}^{t}h(\varphi_{u}(x))\,du\Big),\,Z_{t}(x)=\exp\!\Big(-\int_{0}^{t}\gamma(\varphi_{u}(x))\,du\Big)\,y(x),$
and define the population-weighted effective aggregate process $\eta_{t}$ and the aggregate Duesenberry loading $\partial_{t}\eta_{t}$ as defined in Theorem 3.
Example 2.
Assume the conditions at the beginning of this section (Section 5). Moreover, let $B_{t}(x)$ be a bounded-variation, absolutely continuous process in $t$ with continuous derivative $\partial_{t}B_{t}(x)$ with:
$dB_{t}(x)=\partial_{t}B_{t}(x)\,dt,\qquad B_{t}(x)\geq 0,\qquad\partial_{t}B_{t}(x)\leq 0\qquad\text{for all }x\in\mathbb{D},$
with
$\int_{\mathbb{D}}B_{0}(x)y(x)\,d\mu(x)<\int_{\mathbb{D}}y(x)\,d\mu(x)$ (54)
Also, define the pricing kernel $H_{t}>0$, by Equation (42), where
$H_{t}\coloneqq\frac{-\partial_{t}\eta_{t}}{I_{t}^{\mu}},\qquad\frac{dH_{t}}{H_{t}}=-r_{t}\,dt-\vartheta^{\intercal}_{t}\,dW_{t},\qquad\vartheta_{t}\in\mathbb{R}^{n}.$
Define
$L_{t}(x)=\frac{\eta_{t}B_{t}(x)I^{\mu}_{t}}{\partial_{t}\eta_{t}}\frac{y(x)}{\int_{\mathbb{D}}y(x)\,d\mu(x)},$
where $L_{t}(x)\leq 0$ for all $x\in\mathbb{D}$, and $0\leq t<\infty$. Define,
$Q_{t}(x)=\frac{y(x)}{\int_{\mathbb{D}}y(x)\,d\mu(x)}\frac{I_{t}^{\mu}}{(\partial_{t}\eta_{t})}\partial_{t}\Big(\eta_{t}B_{t}(x)\Big)=\\ \frac{y(x)I_{t}^{\mu}}{(-\partial_{t}\eta_{t})\,\int_{\mathbb{D}}y(x)\,d\mu(x)}\Big(B_{t}(x)(-\partial_{t}\eta_{t})-\eta_{t}\partial_{t}B_{t}(x)\Big)\geq 0.$ (55)
A straightforward computation shows under Assumption 2, that for all $x\in\mathbb{D}$
$\big|H_{t}L_{t}(x)\big|\leq B_{0}(x)y(x),\qquad\int_{0}^{\infty}H_{t}Q_{t}(x)\,dt=y(x)B_{0}(x)$
implying that $L_{t}(x)$ is a subsistence random field and $Q_{t}(x)$ is a rate of endowment structure as in Definition 5. Moreover, $H_{t}L_{t}(x)-\int_{0}^{t}H_{s}Q_{s}(x)\,ds$ is constant and therefore a martingale.
Also, define
$P_{t}^{\mu}=L_{t}^{\mu}+\frac{\eta_{t}}{H_{t}}=\frac{\eta_{t}I^{\mu}_{t}}{\partial_{t}\eta_{t}}\frac{\int_{\mathbb{D}}B_{t}(x)y(x)\,d\mu(x)}{\int_{\mathbb{D}}y(x)\,d\mu(x)}-\frac{\eta_{t}I_{t}^{\mu}}{\partial_{t}\eta_{t}}>0$
for all $t\geq 0$. If we define
$\Sigma_{t}^{I}\coloneqq\int_{\mathbb{D}}\Lambda_{t}(x)\,(\nabla I\,\varrho)(\varphi_{t}(x))\,d\mu(x),\qquad\Sigma_{t}^{\gamma}\coloneqq\int_{\mathbb{D}}Z_{t}(x)\,(\nabla\gamma\,\varrho)(\varphi_{t}(x))\,d\mu(x),$ (56)
by Itô’s formula it follows that the volatility $\sigma_{t}$ of the price process $P^{\mu}_{t}$ and the volatility $\vartheta_{t}$ of the state price process $H_{t}$ are identical:
$\sigma_{t}=\vartheta_{t}=\frac{\Sigma_{t}^{I}}{I_{t}^{\mu}}+\frac{\Sigma_{t}^{\gamma}}{\partial_{t}\eta_{t}}.$ (57)
If $\sigma_{t}$ is non-degenerate for all $t$, then the market satisfies the Equation (35), and the existence construction of Theorem 3 applies. A simple condition which guarantees the non-degenerate condition for $n\geq 2$ is the existence of nonzero vectors $u,v\in\mathbb{R}^{n}$ such that, almost surely with
$u^{\top}\Sigma_{t}^{\gamma}=0,\qquad v^{\top}\Sigma_{t}^{I}=0,\qquad u^{\top}\Sigma_{t}^{I}+v^{\top}\Sigma_{t}^{\gamma}\neq 0,$ (58)
which using Equation (56), which can be interpreted as requiring that income-risk and impatience-risk enter (at least partially) through distinct Brownian directions, so that the two components cannot cancel identically in (57).
Remark 11.
A particular subcase of Example 2, is an economy with no labor income, i.e., $Q_{t}(x)\equiv 0$ for all $(t,x)$ (i.e., $B_{t}(x)\equiv 0$, which also forces $L_{t}(x)\equiv 0$). Then, the total income is entirely generated by the dividend/endowment stream, so the aggregate income reduces to $I_{t}^{\mu}$. The state-price density and aggregate price simplify to
$H_{t}=\frac{-\partial_{t}\eta_{t}}{I_{t}^{\mu}},\qquad P_{t}^{\mu}=\frac{\eta_{t}}{H_{t}}=I_{t}^{\mu}\,\frac{\eta_{t}}{(-\partial_{t}\eta_{t})}.$
Example 3.
Assume the conditions at the beginning of this section (Section 5). Moreover, assume a non-negative continuous random field $\chi_{t}(x)$ such that for each $x\in\mathbb{D}$ the map $t\mapsto\chi_{t}(x)$ is a.s. absolutely continuous, with $\chi_{t}(x)\geq 0$ and $\partial_{t}\chi_{t}(x)\leq 0$ for $t\geq 0,\ x\in\mathbb{D}$. Assume that
$u_{t}(x):=\chi_{t}(x)+\frac{\eta_{t}}{\partial_{t}\eta_{t}}\,\partial_{t}\chi_{t}(x),\qquad\int_{\mathbb{D}}y(x)\,u_{t}(x)\,d\mu(x)<1,\qquad t\geq 0$ (59)
and define
$Q_{t}(x):=y(x)\,I_{t}^{\mu}\,u_{t}(x),\qquad L_{t}(x):=\frac{\eta_{t}}{\partial_{t}\eta_{t}}\,I_{t}^{\mu}\,y(x)\,\chi_{t}(x).$
Since $\partial_{t}\eta_{t}<0$ and $\partial_{t}\chi_{t}(x)\leq 0$, we have $u_{t}(x)\geq\chi_{t}(x)\geq 0$, hence $Q_{t}(x)\geq 0$, while $L_{t}(x)\leq 0$.
Also, Assumption 2 implies that for all $x\in\mathbb{D}$
$\big|H_{t}L_{t}(x)\big|\leq\eta_{0}y(x)\chi_{0}(x),\qquad\int_{0}^{\infty}H_{t}Q_{t}(x)\,dt=\eta_{0}y(x)\chi_{0}(x),$
implying that $L_{t}(x)$ is a subsistence random field and $Q_{t}(x)$ is a rate of endowment as in Definition 5. Moreover, $H_{t}L_{t}(x)-\int_{0}^{t}H_{s}Q_{s}(x)\,ds$ is a constant and therefore a martingale. Define the aggregate price process
$P_{t}^{\mu}:=L_{t}^{\mu}+\frac{\eta_{t}}{H_{t}}=\frac{\eta_{t}I_{t}^{\mu}}{-\partial_{t}\eta_{t}}\Big(1-\int_{\mathbb{D}}y(x)\,\chi_{t}(x)\,d\mu(x)\Big).$
Because $u_{t}(x)\geq\chi_{t}(x)$, (59) implies $\int_{\mathbb{D}}y(x)\chi_{t}(x)\,d\mu(x)<1$, hence $P_{t}^{\mu}>0$. Moreover, since $\chi$ is of finite variation (no Brownian term), the diffusion coefficient of $P^{\mu}$ coincides with that computed in Example 2; in particular the smooth market condition (35) holds (one may take the scalar multiplier to be equal to $1$). Therefore, by Theorem 3, the implied economy is a short-horizon Duesenberry equilibrium.
Alternatively, fix a continuous function $f:\mathbb{D}\to[0,\infty)$ and assume a non-negative random field $u_{t}(x)$ such that for each $x$ the map $t\mapsto u_{t}(x)$ is a.s. absolutely continuous and
$\int_{\mathbb{D}}y(x)\,u_{t}(x)\,d\mu(x)<1,\qquad t\geq 0,$
together with the two inequalities
$\eta_{t}u_{t}(x)+\int_{0}^{t}(-\partial_{s}\eta_{s})u_{s}(x)\,ds=\eta_{0}u_{0}(x)+\int_{0}^{t}\eta_{s}\,\partial_{s}u_{s}(x)\,ds\geq\\ \eta_{0}f(x)\geq\int_{0}^{t}(-\partial_{s}\eta_{s})\,u_{s}(x)\,ds,\qquad t\geq 0,\ x\in\mathbb{D},$ (60)
(For instance, the left-hand side of Equation (60) holds whenever $u_{t}(x)$ is non-decreasing in $t$ and $u_{0}(x)\geq f(x)$.) Define $\chi_{t}(x)$ by the relation
$\chi_{t}(x){\eta_{t}}:={\eta_{0}f(x)+\int_{0}^{t}(\partial_{s}\eta_{s})\,u_{s}(x)\,ds}.$
Then the right-hand side of (60) yields $\chi_{t}(x)\geq 0$, and the left-hand side of Equation (60) is equivalent (via integration by parts) to $\partial_{t}\chi_{t}(x)\leq 0$; moreover,
$u_{t}(x)=\chi_{t}(x)+\frac{\eta_{t}}{\partial_{t}\eta_{t}}\,\partial_{t}\chi_{t}(x).$
Hence, defining
$Q_{t}(x):=y(x)\,I_{t}^{\mu}\,u_{t}(x),\qquad L_{t}(x):=\frac{I_{t}^{\mu}}{\partial_{t}\eta_{t}}\,\,y(x)\,\Big(\eta_{0}f(x)+\int_{0}^{t}(\partial_{s}\eta_{s})\,u_{s}(x)\,ds\Big)$
we have that $H_{t}L_{t}(x)-\int_{0}^{t}H_{s}Q_{s}(x)\,ds$ is constant, therefore a martingale, and with
$P_{t}^{\mu}:=L_{t}^{\mu}+\frac{\eta_{t}}{H_{t}}=\frac{\eta_{t}I_{t}^{\mu}}{-\partial_{t}\eta_{t}}\Big(1-\int_{\mathbb{D}}y(x)\,\chi_{t}(x)\,d\mu(x)\Big)>0.$
As in Example 2 the volatility $\sigma_{t}$ of the price process $P^{\mu}_{t}$ and the volatility of the state price process $\vartheta_{t}$ are given by Equation (57). It also applies that the model satisfies the Smooth Market Condition as long as the condition given by Equation (58) holds. It follows that Theorem 3 yields again, a short-horizon Duesenberry equilibrium.
6. Concluding Remarks
In this paper, we develop a framework for modeling financial markets that can account for the changing behavior of utilities for consumption and investment within infinite populations. We establish explicit sufficient conditions (see Theorem 3) on the primitives of the economy to guarantee the existence of a short-horizon Duesenberry equilibrium without assuming market completeness a priori.
The analysis concludes that incorporating the evolution of consumer preferences regarding relative income offers a robust theoretical framework to address both the “equity premium puzzle” and the “risk-free rate puzzle.”
Furthermore, this equilibrium framework circumvents common theoretical limitations of classical equilibrium theories and provides a tractable method for pricing assets in incomplete markets driven by Brownian flows.
Three directions for future research are particularly promising. First, the interest-rate decomposition (53) and its connection to the consumption–wealth ratio suggest a natural extension to the term structure of interest rates: the Duesenberry loading $\partial_{t}\eta_{t}$ encodes cross-sectional impatience information that should propagate into the yield curve, potentially explaining both the level and the slope of the term structure through demographic fundamentals rather than latent factors. Second, the present framework assumes aggregate income and dividends as exogenous primitives; embedding the model in a production economy, where these processes arise endogenously from production technology and labor supply decisions, would provide a fully general-equilibrium foundation and sharpen the empirical predictions regarding the co-movement of consumption, investment and asset returns. Third, the credit-conditions discussion in Remark 8 points toward Duesenberry equilibria with endogenous borrowing premia, in which the present value of labor income is discounted at a state-dependent household borrowing rate—reflecting default and liquidity risks—rather than at the risk-free rate. This extension arises naturally from combining the equilibrium framework developed here with the life-insurance completeness and state-dependent hazard-rate machinery of Londoño [47]: the latter shows that introducing a state-dependent effective borrowing rate does not alter the aggregate pricing kernel, while the former provides the equilibrium structure in which such rates generate large fluctuations in total wealth through state-dependent discounting of future labor flows. In particular, this approach has the potential to account for the high levels of aggregate wealth volatility implied by the equilibrium (Table 1), where the market price of risk $\vartheta=\sigma^{W}$ consistently exceeds equity volatility $\sigma^{\Sigma}$, thereby offering a complete quantitative resolution of the equity premium puzzle. We develop this program in Londoño [48].
Appendix A Aggregation and Semimartingale Properties
This appendix establishes the semimartingale properties of population-weighted aggregates that underpin the equilibrium analysis in the main text. We first present a general result on aggregated Itô processes (Proposition 1), then develop the cocycle structure of population weights (Lemma 1), and finally specialize to discrete and continuous populations (Propositions 2 and 3).
A.1. Semimartingale property of aggregated processes
In this subsection we discuss some properties of the aggregation of processes of the form $\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)$ where $\varphi$ is a consistent semimartingale process, and $\mu$ is a probability distribution on $\mathbb{D}$. Such processes arise naturally in the context of modeling infinite populations, where each individual is represented by a point $x\in\mathbb{D}$, and $\varphi_{s,t}(x)$ represents the stochastic behavior of the individual $x$ at time $t$ given that at time $s$ the state of this individual was $x$. The function $f(t,x)$ represents the relationship between the state of the individual at time $t$ and some variable of interest. Finally, the measure $\mu$ represents the initial distribution of individuals at time $s$. In this section, we provide conditions under which such aggregated processes are semimartingales and provide their stochastic representations.
Proposition 1.
Let $f\colon[0,\infty)\times\mathbb{D}\to\mathbb{R}$ be a continuous function with continuous partial derivatives $\partial^{2}f/\partial x_{i}\partial x_{j}$ and $\partial f/\partial t$, for $i,j=1\cdots d$ ($f\in C^{1,2}([0,\infty)\times\mathbb{D}$), where $\mathbb{D}\subset\,\mathbb{R}^{d}$ is an open subset. Suppose that $\varphi$ is consistent semimartingale process with values in $\mathbb{D}$ with stochastic representation (2) where $\rho$ and $\varrho$ are functions of class $C^{m,\chi}$ where $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$. Assume that $\mu$ is a probability distribution on $\mathbb{D}$ with a continuous density $g(x)$. Assume that for any $s<t<\infty$:
  1. i
    $f(t,\cdot)$ is $\mu$ integrable for any $t$.
  2. $i$
    $\partial f/\partial t$, $\sum_{i}\rho^{i}\partial f/\partial x_{i}$, $\sum_{i}\varrho^{i,k}\partial f/\partial x^{i}$, $\sum_{i,j}\varrho^{i,j}\partial^{2}f/\partial x_{i}\partial x_{j}$ are Lipschitz continuous for all $k$, where $\varrho^{k}=(\varrho^{1,k},\cdots,\varrho^{d,k})$ and $\rho=(\rho^{1},\cdots,\rho^{d})$ for $k=1,\cdots,n$.
  3. iii
    $E[\int_{\mathbb{D}}|\varphi^{i}_{s,t}(x)|d\mu(x)]<\infty$ for $i=1,\cdots,d$.
Then, for each fixed $s$ the process $\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)$ is a continuous $(\mathcal{F}^{T}_{s,t})_{t\in[s,T]}$ semimartingale process almost $P$-a.s, and
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\mathbb{D}}f(s,x)g(x)\,dx+\int_{s}^{t}\int_{\mathbb{D}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x))g(x)\,dx\,du+\\ \int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x))\rho^{i}(u,\varphi_{s,u}(x))g(x)dx\,du+\\ \frac{1}{2}\int_{s}^{t}\sum_{i,j}\int_{\mathbb{D}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x))\varrho^{i,j}(u,\varphi_{s,u}(x)\right)g(x)\,dx\,du+\\ \sum_{k}\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x))\varrho^{i,k}(u,\varphi_{s,u}(x))g(x)\,dx\,dW_{s}^{k}(u).$ (61)
Proof.
We introduce some notation. Let $a=(a_{1},\cdots,a_{d})<b=(b_{1},\cdots,b_{d})$, and let $P_{i}=\{t^{i}_{0}=a_{i}<t^{i}_{1}<,\cdots,<t^{i}_{p_{i}}=b_{i}\}$ for $i=1\cdots d$ be a partition of the hyper-rectangle $[a,b]$. For each $t=(t^{1}_{i_{1}},\cdots,t^{d}_{i_{d}})\in P=P_{1}\times\cdots\times P_{d}$, with $t<b$ we define the next corner $t^{+}=(t^{1}_{i_{1}+1},\cdots,t^{d}_{i_{d}+1})$. We also define a sample for each $t\in P$ to be a point $t^{\star}\in[t,t^{+})$, where for $t,t^{+})=\{x\in\mathbb{R}^{d}\colon t\leq x<t^{+}\}$. To each partition $P$, we define the norm or mesh
$|P|=\max_{t\in P,t<b}{\|t-t^{+}\|}$
where $\|x\|=\sqrt{x^{\intercal}x}$ denotes the $L^{2}$ norm of the vector $x\in\mathbb{R}^{d}$.
Let $U\subset\mathbb{D}$ be an open set with compact closure $\bar{U}\subset[a,b]$, and assume that $f(t,x)=0$ for $x\notin U$. It follows from Itô’s rule that if $(P_{r})$ is a sequence of partitions of a hyper-rectangle $[a,b]$, then if we denote $m(\cdot)$ as the Lebesgue measure on $\mathbb{D}$,
$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\bar{U}}f(t,\varphi_{s,t}(x))g(x)\,dm=\lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,\varphi_{s,t}(x^{\star}))g(x^{\star})m([x,x^{+}))=\\ \lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,x^{\star})g(x^{\star})m([x,x^{+}))+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{x\in P_{r}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x^{\star}))\rho^{i}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\frac{1}{2}\int_{s}^{t}\sum_{i,j}\sum_{x\in P_{r}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x^{\star}))\varrho^{i,j}(u,\varphi_{s,u}(x^{\star})\right)g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\sum_{k}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x^{\star}))\varrho^{i,k}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,dW_{s}^{k}(u),$ (62)
where the validity of Equation (61) is a consequence of the classical dominated convergence theorem (where we use the dominated convergence theorem for stochastic integrals for the last term). We notice that we used the fact that all integrands are 0 outside a compact set in the last four terms. For the general case, let ${U}_{r}$ be a sequence of open sets with compact closure, with $\bar{U}_{r}\subset U_{r+1}$ and $\cup_{r}U_{r}=\mathbb{D}$ and let $h_{r}(x)$ be a sequence of $C^{\infty}$ functions such that $h_{r}(x)=1$ on $\bar{U}_{r}$ and $h_{r}(x)=0$ on $U_{r+1}$. Then $f_{r}(t,x)=h_{r}(x)f(t,x)$ is a sequence of functions with the same degree of smoothness with $f_{r}(t,x)\to f(t,x)$ as $r\to\infty$ for all $(t,x)\in[0,\infty)\times\mathbb{D}$. Consequently, based on the statements proved above and the Equation (61), the proposition holds for each $f_{r}$. Finally, we obtain the result using the dominated convergence theorem (classical and stochastic versions). We note that conditions (i)–(iii) imply that a continuous process dominates the terms of the integrals. ∎
Remark 12.
If $\rho$ and $\varrho$ are uniformly Lipschitz continuous defined on $\mathbb{R}^{d}$ then there exists $K>0$ such that $|E[\varphi_{s,t}(x)-x]|\leq K(1+|x|(t-s))$ for $0\leq s\leq t$ (see Kunita [37, Theorem 4.2.5]). It follows that under the assumption of Lipschitz continuity for $\rho$ and $\varrho$, assuming $\int_{\mathbb{D}}|x_{i}|g(x)\,dm<\infty$ for all $i$ where $d\mu/dm=g$, implies that condition iiiiii of Proposition 1 holds. The latter follows using the dominated convergence theorem for stochastic integrals.
A.2. Cocycle properties of population weights
Lemma 1 (Cocycle and identity properties).
Adopt Definition 2. The following cocycle and identity properties hold $P$-a.s.:
  1. (i)
    (Multiplicative weight cocycle.) For all $0\leq r\leq s\leq t$ and $x\in\mathbb{D}$,
    $\Lambda_{r,t}(x)=\Lambda_{r,s}(x)\,\Lambda_{s,t}(\varphi_{r,s}(x)).$ (63)
  2. (ii)
    (Kernel cocycle.) For all $0\leq r\leq s\leq t$ and Borel $B\subset\mathbb{D}$,
    $\nu_{r,t}(x,B)=\int_{\mathbb{D}}\nu_{s,t}(y,B)\,\nu_{r,s}(x,dy),\qquad\text{i.e.}\qquad\nu_{r,t}=\nu_{r,s}\circ\nu_{s,t}.$ (64)
  3. (iii)
    (Measure cocycle.) For any (deterministic or $\mathcal{F}_{r}$-measurable) finite measure $\mu$ on $\mathbb{D}$,
    $\mu_{r,t}=(\mu_{r,s})\nu_{s,t}.$ (65)
  4. (iv)
    (0-based aggregate identity.) Fix $\mu$ at time 0 and set $\mu_{s}\coloneqq\mu_{0,s}$ as in Definition 2. Then for any Borel function $f$ such that the integrals below are finite, and any $0\leq s\leq t$,
    $\int_{\mathbb{D}}f(\varphi_{s,t}(x))\,\Lambda_{s,t}(x)\,\mu_{s}(dx)=\int_{\mathbb{D}}f(\varphi_{0,t}(y))\,\Lambda_{0,t}(y)\,\mu(dy).$ (66)
    Equivalently, the population-weighted aggregate computed on $[s,t]$ from the shifted population $\mu_{s}$ coincides pathwise with the aggregate computed on $[0,t]$ from the initial population $\mu$.
Proof.
Identity (63) follows from additivity of the time integral and the flow property $\varphi_{r,u}=\varphi_{s,u}\circ\varphi_{r,s}$ for $u\geq s$. Then (64) is immediate since
$\nu_{r,t}(x,B)=\Lambda_{r,t}(x)\mathbf{1}_{\{\varphi_{r,t}(x)\in B\}}=\Lambda_{r,s}(x)\Lambda_{s,t}(\varphi_{r,s}(x))\mathbf{1}_{\{\varphi_{s,t}(\varphi_{r,s}(x))\in B\}},$
and $\nu_{r,s}(x,\cdot)=\Lambda_{r,s}(x)\delta_{\varphi_{r,s}(x)}(\cdot)$.
Identity (65) is obtained by integrating (64) against $\mu$.
Finally, (66) is just (65) with $(r,s,t)=(0,s,t)$ written in test-function form:
$\int f(z)\,\mu_{t}(dz)=\int f(\varphi_{s,t}(x))\Lambda_{s,t}(x)\,\mu_{s}(dx),$
and $\int f(z)\,\mu_{t}(dz)=\int f(\varphi_{0,t}(y))\Lambda_{0,t}(y)\,\mu(dy)$ by definition of $\mu_{t}$. ∎
A.3. Discrete populations
Assumption 3 (Discrete population and population weights).
Assume a diffeomorphic Brownian flow that satisfies Assumption 1. Fix $s\geq 0$. Let the population be discrete,
$\mu=\sum_{i\in I}w_{i}\,\delta_{x_{i}},\qquad\sum_{i\in I}|w_{i}|<\infty,$
where $I$ is finite or countable and $x_{i}\in\mathbb{D}$. Let $h:\mathbb{D}\to\mathbb{R}$ be a continuous $C^{2,0+}$ function and define the population weight as in Definition 2, where it is assumed that $\mathcal{P}_{\varphi}$ is the set of finite measures. Assume that for each $i\in I$, $t\mapsto\psi^{\mu}_{t}(x_{i})$ is a continuous semimartingale with decomposition
$\psi^{\mu}_{t}(x_{i})=\psi^{\mu}_{0}(x_{i})+\int_{0}^{t}\zeta_{u}^{\mu}a\big(\varphi_{u}(x_{i})\big)\,du+\int_{0}^{t}(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x_{i}))\big)\,dW(u),$
for continuous functions $a$ (real-valued) and $b$ ($\mathbb{R}^{n}$-valued), consistent with respect to the population structure $\boldsymbol{\mu}$, and continuous semimartingale processes real valued $\zeta_{u}^{\mu}$, and $\mathbb{R}^{n}$ valued $(\upsilon^{\mu})_{u}$ (which is type independent). Assume either $I$ is finite, or for every $T>0$,
$\sum_{i\in I}|w_{i}|\int_{0}^{T}\Big(\big|\Lambda_{u}(x_{i})\,\zeta_{u}^{\mu}a\big(\varphi_{u}(x_{i})\big)\big|+\big|\Lambda_{u}(x_{i})\,h\big(\varphi_{u}(x_{i})\big)\,\psi_{u}(x_{i})\big|\Big)\,du<\infty,$
and
$\sum_{i\in I}|w_{i}|\,\Big(\int_{0}^{T}\Lambda_{u}(x_{i})^{2}\,\big\|(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x_{i}))\big)\big\|^{2}\,du\Big)^{1/2}<\infty.$
Proposition 2 (Population-weighted aggregation: discrete population).
Under Assumption 3, define the population-weighted aggregate (cf. Definition 2)
$\psi^{\mu}_{t}\coloneqq\int_{\mathbb{D}}\Lambda_{t}(x)\,\psi^{\mu}_{t}(x)\,d\mu(x)=\sum_{i\in I}w_{i}\,\Lambda_{t}(x_{i})\,\psi_{t}(x_{i}).$
Then, $t\mapsto\psi^{\mu}_{t}$ is a continuous semimartingale for $0\leq t<\infty$.
Moreover,
$\psi^{\mu}_{t}$ $=\psi^{\mu}_{0}+\int_{0}^{t}\sum_{i\in I}w_{i}\Big[\Lambda_{u}(x_{i})\,\zeta_{u}^{\mu}a\big(\varphi_{u}(x_{i})\big)+\Lambda_{u}(x_{i})\,h\big(\varphi_{u}(x_{i})\big)\,\psi_{u}(x_{i})\Big]\,du$
$\quad+\int_{0}^{t}\sum_{i\in I}w_{i}\,\Lambda_{u}(x_{i})\,(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x_{i}))\big)\,dW(u).$
Proof.
For each $i$, $\Lambda_{t}(x_{i})$ has finite variation with $d\Lambda_{t}(x_{i})=\Lambda_{t}(x_{i})\,h(\varphi_{t}(x_{i}))\,dt$. Hence, $\Lambda_{t}(x_{i})\psi_{t}(x_{i})$ is a continuous semimartingale by integration by parts, with drift and diffusion as stated. The summability conditions justify exchanging the (countable) sum with the Lebesgue and stochastic integrals, which yield the aggregate decomposition. ∎
A.4. Continuous populations
Assumption 4 (Continuous populations).
Assume a diffeomorphic Brownian flow that satisfies Assumption 1. Let $h:\mathbb{D}\to\mathbb{R}$ be a continuous $C^{2,0+}$ function and define the population weight as in Definition 2. Let $\mu$ be a finite Borel measure. Assume that for each $x\in\mathbb{D}$,
$\psi^{\mu}_{t}(x)=\psi_{0}(x)+\int_{0}^{t}\zeta_{u}^{\mu}a\big(\varphi_{u}(x)\big)\,du+\int_{0}^{t}(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x))\big)\,dW(u)$
is a continuous semimartingale in $t$, for continuous real-valued functions $a$ and $\mathbb{R}^{n}$-valued $b$, consistent with respect to the population structure $\boldsymbol{\mu}$, and continuous semimartingale processes real valued $\zeta_{u}^{\mu}$, and $\mathbb{R}^{n}$ valued $(\upsilon^{\mu})_{u}$ (which is type independent).
Assume that for every $T>0$, almost everywhere
$\int_{\mathbb{D}}\int_{0}^{T}\Big(\big|\Lambda_{u}(x)\,\zeta_{u}^{\mu}a\big(\varphi_{u}(x)\big)\big|+\big|\Lambda_{u}(x)\,h\big(\varphi_{u}(x)\big)\,\psi_{u}(x)\big|\Big)\,du\,d\mu(x)<\infty,$
and
$\int_{\mathbb{D}}\Big(\int_{0}^{T}\Lambda_{u}(x)^{2}\,\big\|(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x))\big)\big\|^{2}\,du\Big)^{1/2}\,d\mu(x)<\infty.$
Proposition 3 (Aggregate semimartingale: continuous population).
Under Assumption 4, the population-weighted aggregate
$\psi^{\mu}_{t}\coloneqq\int_{\mathbb{D}}\psi_{t}(x)\,\Lambda_{t}(x)\,d\mu(x)$
is a continuous semimartingale for $0\leq t<\infty$.
Moreover,
$\psi^{\mu}_{t}$ $=\psi^{\mu}_{0}+\int_{0}^{t}\int_{\mathbb{D}}\Big[\Lambda_{u}(x)\,\zeta_{u}^{\mu}a\big(\varphi_{u}(x)\big)+\Lambda_{u}(x)\,h\big(\varphi_{u}(x)\big)\,\psi_{u}(x)\Big]\,d\mu(x)\,du$
$\quad+\int_{0}^{t}\int_{\mathbb{D}}\Lambda_{u}(x)\,(\upsilon_{u}^{\mu})^{\intercal}\mathrm{diag}\big(b(\varphi_{u}(x))\big)\,d\mu(x)\,dW(u).$
Proof.
Apply the product rule to $x\mapsto\psi_{t}(x)\Lambda_{t}(x)$ to identify its drift and diffusion from Assumption 4. Then integrate in $x$ and justify exchanging the $x$-integral with the Lebesgue/stochastic integrals by the same dominated-convergence and stochastic-Fubini argument used in Proposition 1, proving the stated decomposition. ∎
Corollary 4 (Aggregate semimartingale and restart property).
Assume Assumption 3 or Assumption 4, and adopt the population structure of Lemma 1. Fix $t\geq 0$ and set
$\mu_{t}\coloneqq\mu_{0,t}.$
(Restarted aggregate process). For $u\geq 0$, define the restarted aggregate by
$\psi^{\mu_{t}}_{0,u}\coloneqq\int_{\mathbb{D}}\bar{\psi}_{t,t+u}(x)\,\Lambda_{t,t+u}(x)\,d\mu_{t}(x)\;=\;\int_{\mathbb{D}}\psi_{t,t+u}(x)\,d\mu_{t}(x).$
Then $u\mapsto\psi^{\mu_{t}}_{0,u}$ is a continuous semimartingale on $0,\infty)$.
Moreover, its drift and diffusion coefficients are the corresponding population-weighted aggregates of the individual coefficients computed with respect to the shifted initial population $\mu_{t}$ and the shifted flow $\varphi_{t,t+u}$, $u\geq 0$.
(Pathwise identification with calendar time). If $\psi$ is $\varphi$-consistent, then for every $u\geq 0$,
$\psi^{\mu_{t}}_{0,u}\;=\;\psi^{\mu}_{0,t+u}\;=:\;\psi^{\mu}_{t+u}.$
In particular, the restarted process $u\mapsto\psi^{\mu_{t}}_{0,u}$ coincides pathwise with the calendar-time process $u\mapsto\psi^{\mu}_{t+u}$.

### 18 Mar 2026 v2

@@
 Proposition 1.
-Let $f\colon[0,\infty)\times\mathbb{D}\to\mathbb{R}$ be a continuous function with continuous partial derivatives $\partial^{2}f/\partial x_{i}\partial x_{j}$ and $\partial f/\partial t$, for $i,j=1\cdots d$ ($f\in C^{1,2}([0,\infty)\times\mathbb{D}$), where $\mathbb{D}\subset\,\mathbb{R}^{d}$ is an open subset.
-Suppose that $\varphi$ is consistent semimartingale process with values in $\mathbb{D}$ with stochastic representation (2) where $\rho$ and $\varrho$ are functions of class $C^{m,\chi}$ where $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$.
-Assume that $\mu$ is a probability distribution on $\mathbb{D}$ with a continuous density $g(x)$.
+Let $f\colon[0,\infty)\times\mathbb{D}\to\mathbb{R}$ be a continuous function with continuous partial derivatives $\partial^{2}f/\partial x_{i}\partial x_{j}$ and $\partial f/\partial t$, for $i,j=1,\dots,d$ (that is, $f\in C^{1,2}([0,\infty)\times\mathbb{D})$), where $\mathbb{D}\subset\mathbb{R}^{d}$ is an open subset.
+Suppose that $\varphi$ is a consistent semimartingale process with values in $\mathbb{D}$ and stochastic representation (2), where $\rho$ and $\varrho$ are functions of class $C^{m,\chi}$ with $m=0$ and $\chi=1$ or $m\geq 1$, and $\chi>0$.
+Assume that $\mu$ is a probability distribution on $\mathbb{D}$ with continuous density $g(x)$.
+Define
+$a^{ij}(t,x)\coloneqq\sum_{k=1}^{n}\varrho^{i,k}(t,x)\varrho^{j,k}(t,x),\qquad i,j=1,\dots,d.$
 Assume that for any $s<t<\infty$:
   1. i
-    $f(t,\cdot)$ is $\mu$ integrable for any $t$.
+    $f(t,\cdot)$ is $\mu$-integrable for any $t$.
   2. $i$
-    $\partial f/\partial t$, $\sum_{i}\rho^{i}\partial f/\partial x_{i}$, $\sum_{i}\varrho^{i,k}\partial f/\partial x^{i}$, $\sum_{i,j}\varrho^{i,j}\partial^{2}f/\partial x_{i}\partial x_{j}$ are Lipschitz continuous for all $k$, where $\varrho^{k}=(\varrho^{1,k},\cdots,\varrho^{d,k})$ and $\rho=(\rho^{1},\cdots,\rho^{d})$ for $k=1,\cdots,n$.
+    $\partial f/\partial t$, $\sum_{i}\rho^{i}\,\partial f/\partial x_{i}$, $\sum_{i}\varrho^{i,k}\,\partial f/\partial x_{i}$ for every $k=1,\dots,n$, and $\sum_{i,j}a^{ij}\,\partial^{2}f/\partial x_{i}\partial x_{j}$ are Lipschitz continuous.
   3. iii
-    $E[\int_{\mathbb{D}}|\varphi^{i}_{s,t}(x)|d\mu(x)]<\infty$ for $i=1,\cdots,d$.
+    $E\!\left[\int_{\mathbb{D}}|\varphi^{i}_{s,t}(x)|\,d\mu(x)\right]<\infty$ for $i=1,\dots,d$.
 Then, for each fixed $s$ the process $\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)$ is a continuous $(\mathcal{F}^{T}_{s,t})_{t\in[s,T]}$ semimartingale process almost $P$-a.s, and
-$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\mathbb{D}}f(s,x)g(x)\,dx+\int_{s}^{t}\int_{\mathbb{D}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x))g(x)\,dx\,du+\\ \int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x))\rho^{i}(u,\varphi_{s,u}(x))g(x)dx\,du+\\ \frac{1}{2}\int_{s}^{t}\sum_{i,j}\int_{\mathbb{D}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x))\varrho^{i,j}(u,\varphi_{s,u}(x)\right)g(x)\,dx\,du+\\ \sum_{k}\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x))\varrho^{i,k}(u,\varphi_{s,u}(x))g(x)\,dx\,dW_{s}^{k}(u).$ (61)
+$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\int_{\mathbb{D}}f(s,x)g(x)\,dx\\ +\int_{s}^{t}\int_{\mathbb{D}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x))g(x)\,dx\,du\\ +\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x))\rho^{i}(\varphi_{s,u}(x))g(x)\,dx\,du\\ +\frac{1}{2}\int_{s}^{t}\sum_{i,j}\int_{\mathbb{D}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f(u,\varphi_{s,u}(x))a^{ij}(\varphi_{s,u}(x))g(x)\,dx\,du\\ +\sum_{k}\int_{s}^{t}\sum_{i}\int_{\mathbb{D}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x))\varrho^{i,k}(\varphi_{s,u}(x))g(x)\,dx\,dW_{s}^{k}(u).$ (61)
 Proof.
@@
 It follows from Itô’s rule that if $(P_{r})$ is a sequence of partitions of a hyper-rectangle $[a,b]$, then if we denote $m(\cdot)$ as the Lebesgue measure on $\mathbb{D}$,
-$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\bar{U}}f(t,\varphi_{s,t}(x))g(x)\,dm=\lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,\varphi_{s,t}(x^{\star}))g(x^{\star})m([x,x^{+}))=\\ \lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,x^{\star})g(x^{\star})m([x,x^{+}))+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{x\in P_{r}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x^{\star}))\rho^{i}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\frac{1}{2}\int_{s}^{t}\sum_{i,j}\sum_{x\in P_{r}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\left(u,\varphi_{s,u}(x^{\star}))\varrho^{i,j}(u,\varphi_{s,u}(x^{\star})\right)g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\sum_{k}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x^{\star}))\varrho^{i,k}(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,dW_{s}^{k}(u),$ (62)
+$\int_{\mathbb{D}}f(t,\varphi_{s,t}(x))\,d\mu(x)=\\ \int_{\bar{U}}f(t,\varphi_{s,t}(x))g(x)\,dm=\lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,\varphi_{s,t}(x^{\star}))g(x^{\star})m([x,x^{+}))=\\ \lim_{|P_{r}|\to 0}\sum_{x\in P_{r}}f(t,x^{\star})g(x^{\star})m([x,x^{+}))+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{x\in P_{r}}\frac{\partial}{\partial t}f(u,\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial}{\partial x_{i}}f(u,\varphi_{s,u}(x^{\star}))\rho^{i}(\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\frac{1}{2}\int_{s}^{t}\sum_{i,j}\sum_{x\in P_{r}}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f\bigl(u,\varphi_{s,u}(x^{\star})\bigr)a^{ij}\bigl(\varphi_{s,u}(x^{\star})\bigr)g(x^{\star})m([x,x^{+}))\,du+\\ \lim_{|P_{r}|\to 0}\sum_{k}\int_{s}^{t}\sum_{i}\sum_{x\in P_{r}}\frac{\partial f}{\partial x_{i}}(u,\varphi_{s,u}(x^{\star}))\varrho^{i,k}(\varphi_{s,u}(x^{\star}))g(x^{\star})m([x,x^{+}))\,dW_{s}^{k}(u).$ (62)
 where the validity of Equation (61) is a consequence of the classical dominated convergence theorem (where we use the dominated convergence theorem for stochastic integrals for the last term).